### v1

# Topological rigidity of small RCD(K,N) spaces with maximal rank

Abstract.
For a polycyclic group $\Lambda$, ${\text{rank}}(\Lambda)$ is defined as the number of $\mathbb{Z}$ factors in a polycyclic decomposition of $\Lambda$. For a finitely generated group $G$, ${\text{rank}}(G)$ is defined as the infimum of ${\text{rank}}(\Lambda)$ among finite index polycyclic subgroups $\Lambda\leq G$.
For a compact $\text{RCD}(K,N)$ space $(X,\mathsf{d},\mathfrak{m})$ with ${\text{diam}}(X)\leq\varepsilon(K,N)$, the rank of $\pi_{1}(X)$ is at most $N$. In this note we show that in case of equality, $X$ is homeomorphic to an infranilmanifold, generalizing to the non-smooth setting a result by Kapovitch–Wilking. We also fill a gap in the proof that if $\pi_{1}(X)=\mathbb{Z}^{N}$, then $X$ is bi-Hölder homeomorphic to a flat torus (diffeomorphic in the smooth case).
1. Introduction
A classical result of Gromov [G78], refined by Ruh [R82], asserts that a closed $N$-dimensional Riemannian manifold $X$ with
$\hskip 22.76228pt|\sec(X)|\leq K\hskip 85.35826pt{\text{diam}}(X)\leq\varepsilon(K,N)$
is diffeomorphic to an infranilmanifold. This full topological control is lost if one works with only a Ricci curvature bound; this is illustrated by the fact that the topology of Ricci flat manifolds is still quite mysterious. However, the fundamental group is very well understood. By the work of Kapovitch–Wilking [KW11], if $X$ satisfies
$\hskip 25.6073pt\text{Ric}(X)\geq K\hskip 85.35826pt{\text{diam}}(X)\leq\varepsilon(K,N)$
then $\pi_{1}(X)$ contains a subgroup $\Lambda$ of index $[\pi_{1}(X):\Lambda]\leq C(K,N)$ generated by elements $\{u_{1},\ldots,u_{N}\}\subset\Lambda$ with $[u_{i},u_{j}]\in\langle u_{j+1},\ldots,u_{N}\rangle$ for each $i,j\in\{1,\ldots,N\}$. Moreover, if
$\langle u_{j},\ldots,u_{N}\rangle/\langle u_{j+1},\ldots,u_{N}\rangle\cong\mathbb{Z}\text{ for all }j\in\{1,\ldots,N\},$
then $X$ is homeomorphic to an infranilmanifold. The main result of this note is an extension of this maximal rank topological rigidity to the non-smooth setting.
[…]
For a polycyclic group $\Lambda$, ${\text{rank}}(\Lambda)$ is defined as the number of $\mathbb{Z}$ factors in a polycyclic decomposition of $\Lambda$. For a finitely generated group $G$, ${\text{rank}}(G)$ is defined as the infimum of ${\text{rank}}(\Lambda)$ among finite index polycyclic subgroups $\Lambda\leq G$111we consider the infimum of the empty set to be +$\infty$+ ∞.
[…]
For each $K\in\mathbb{R},$ $N\geq 1$, there is $\varepsilon>0$ such that if $(X,\mathsf{d},\mathfrak{m})$ is an $\text{RCD}(K,N)$ space with ${\text{diam}}(X)\leq\varepsilon$, then ${\text{rank}}(\pi_{1}(X))\leq N$, and in case of equality, $X$ is homeomorphic to an infranilmanifold of dimension $N$.
The Borel conjecture asserts that any two closed aspherical manifolds with isomorphic fundamental group are homeomorphic [R83]. Since this conjecture is known to be true for manifolds with virtually nilpotent fundamental groups (see Theorem 3), and the Margulis Lemma holds for $\text{RCD}(K,N)$ spaces (see [DSZZ23]), then the proof of Theorem 1 consists mainly of showing that if ${\text{diam}}(X)\leq\varepsilon$ and ${\text{rank}}(\pi_{1}(X))=N$, then $X$ is topologically a closed aspherical manifold. This approach is also used by Kapovitch to show that an $\text{RCD}(K,N)$ space with an upper sectional curvature bound, small diameter, and empty boundary is homeomorphic to an infranilmanifold [K21].
1.1. Bi-Hölder homeomorphisms
Theorem 1 can be regarded as a non-abelian analogue of the following result [MMP20, Theorem 1.2.3]. Recall that for a path connected topological space $X$, its first Betti number $\text{b}_{1}(X)$ is defined as the supremum of $k$ for which there is a surjective morphism $\pi_{1}(X)\to\mathbb{Z}^{k}$.
[…]
For each $K\in\mathbb{R},$ $N\geq 1$, there is $\varepsilon>0$ such that if $(X,\mathsf{d},\mathfrak{m})$ is an $\text{RCD}(K,N)$ space with ${\text{diam}}(X)\leq\varepsilon$, then $\text{b}_{1}(X)\leq N$, and in case of equality, $X$ is bi-Hölder homeomorphic to a flat torus of dimension $N$.
The other goal of this note is to fill a small gap in the original proof of the equality case in Theorem 1.1, also present in the proof of the smooth analogue [CC97, Theorem A.1.13]. We now give an overview of this gap. By contradiction, one assumes there is a sequence of $\text{RCD}(K,N)$ spaces $(X_{i},\mathsf{d}_{i},\mathfrak{m}_{i})$ with
${\text{diam}}(X_{i})\to 0\hskip 85.35826pt\text{b}_{1}(X_{i})=N$
but $X_{i}$ not bi-Hölder homeomorphic to a flat torus. In [MMP20], a sequence of finite sheeted covers $X_{i}^{\prime}\to X_{i}$ is constructed in such a way that $X_{i}^{\prime}$ converges in the Gromov–Hausdorff sense to the flat torus $\mathbb{T}^{N}$. From here, the Cheeger–Colding version of Reifenberg theorem [CC97, Appendix 1] implies that $X_{i}^{\prime}$ is bi-Hölder homeomorphic to $\mathbb{T}^{N}$ for $i$ large enough, and it is deduced using topological arguments that the spaces $X_{i}$ are homeomorphic to $\mathbb{T}^{N}$. This concludes their argument, but it doesn’t finish the proof, since a bi-Hölder homeomorphism $X_{i}^{\prime}\to\mathbb{T}^{N}$ doesn’t necessarily descend to a bi-Hölder homeomorphism between $X_{i}$ and a flat torus.
The fix to this gap is pretty clear: if $G_{i}\leq\text{Iso}(X_{i}^{\prime})$ denotes the group of deck transformations, then $X_{i}=X_{i}^{\prime}/G_{i}$, and one needs to find
  • A free isometric action of $G_{i}$ on $\mathbb{T}^{N}$.
  • A $G_{i}$-equivariant bi-Hölder homeomorphism $X_{i}^{\prime}\to\mathbb{T}^{N}$.
This is achieved by the canonical homeomorphisms constructed by Honda–Peng in [HP23], inspired by the work of Wang–Zhao [WZ23], and built upon the techniques and tools developed by Cheeger–Jiang–Naber [CJN21]. We cover the details in Section 5.
Under the conditions of Theorem 1, if $X$ is a smooth $N$-dimensional Riemannian manifold, then it follows from [NZ16, Proposition 5.4] and the work of Huang–Kong–Rong–Xu [HKRX20] that $X$ is diffeomorphic to an infranilmanifold (this was also proven by Naber–Zhang [NZ16] assuming an upper Ricci curvature bound). We note, however, that their proof relies on Ricci flow, so their techniques do not extend to the non-smooth setting. Later Rong [R20] gave a more geometric proof inspired by Cheeger–Fukaya–Gromov collapsing theory [CFG92], but some tools used there still rely on the smooth structure. In view of this discussion and Theorem 1.1, it would be interesting to know if the conclusion of Theorem 1 can be improved to a bi-Hölder homeomorphism.
[…]
For each $K\in\mathbb{R},$ $N\geq 1$, there is $\varepsilon>0$ such that if $(X,\mathsf{d},\mathfrak{m})$ is an $\text{RCD}(K,N)$ space with ${\text{diam}}(X)\leq\varepsilon$, then ${\text{rank}}(\pi_{1}(X))\leq N$, and in case of equality, $X$ is bi-Hölder homeomorphic to an infranilmanifold of dimension $N$.
The following proposition shows that the approach of constructing finite-sheeted covers $X_{i}^{\prime}\to X_{i}$ that don’t collapse fails when dealing with Conjecture 1.1 because in general, fundamental groups of infranilmanifolds are not virtually abelian.
[…]
Let $X_{i}$ be a sequence of closed $N$-dimensional Riemannian manifolds with $\text{Ric}(X_{i})\geq-\frac{1}{i}$ and $\pi_{1}(X_{i})$ torsion-free for all $i$. If there is a sequence of finite sheeted covers $X_{i}^{\prime}\to X_{i}$ with
${\text{diam}}(X_{i}^{\prime})\leq D\hskip 56.9055pt\mathcal{H}^{N}(X_{i}^{\prime})\geq\nu$
for some $D,\nu>0$, then $\pi_{1}(X_{i})$ is virtually abelian for large $i$.
Proof.
Arguing by contradiction, after passing to a subsequence, we can assume $\pi_{1}(X_{i})$ is not virtually abelian for any $i$. Let $\tilde{X}_{i}$ be the sequence of universal covers, $p_{i}\in X_{i}^{\prime}$, and $\tilde{p}_{i}\in\tilde{X}_{i}$ in the preimage of $p_{i}$. After further passing to a subsequence, the triples $(\tilde{X}_{i},\tilde{p}_{i},\pi_{1}(X_{i}^{\prime}))$ converge in the equivariant Gromov–Hausdorff sense to a triple $(Y,\tilde{p},G)$, while $X_{i}^{\prime}$ converges to $Y/G$.
Since the sequence $X_{i}^{\prime}$ doesn’t collapse, neither does the sequence $\tilde{X}_{i}$, and the limit group $G$ is discrete. By Corollary 2.3, the groups $\pi_{1}(X_{i}^{\prime})$ admit no small subgroups hence they are isomorphic to $G$ for large $i$. By [PW22, Theorem 1.4], $\pi_{1}(Y/G)$ is the quotient of $G$ by a subgroup generated by torsion elements, but since $G$ is torsion-free such subgroup must be trivial and $\pi_{1}(Y/G)=G$.
Finally, $Y/G$ equipped with $\mathcal{H}^{N}$ is an $\text{RCD}(0,N)$ space so its fundamental group is virtually abelian, hence so is $\pi_{1}(X_{i})$ for large $i$; a contradiction. ∎
If $\sec(X_{i})\geq-1/i$, one could remove the hypothesis that the groups $\pi_{1}(X_{i})$ are torsion-free since one could use Perelman stability theorem [K07, P91] to conclude that $X_{i}^{\prime}$ and $Y/G$ are homeomorphic for large $i$. The above proposition is illustrated by the following example.
[…]
Consider the $3$-dimensional Heisenberg groups
$H^{3}(\mathbb{R})=\left\{\left.\begin{bmatrix}1&x&z\\ 0&1&y\\ 0&0&1\end{bmatrix}\right|x,y,z\in\mathbb{R}\right\},\quad H^{3}(\mathbb{Z})=\left\{\left.\begin{bmatrix}1&a&c\\ 0&1&b\\ 0&0&1\end{bmatrix}\right|a,b,c\in\mathbb{Z}\right\}.$
Let $\{\text{v}_{1},\,\text{v}_{2},\,\text{v}_{3}\}$ be the basis of the Lie algebra of $H^{3}(\mathbb{R})$ given by
$\text{v}_{1}=\begin{bmatrix}0&1&0\\ 0&0&0\\ 0&0&0\end{bmatrix}\hskip 28.45274pt\text{v}_{2}=\begin{bmatrix}0&0&0\\ 0&0&1\\ 0&0&0\end{bmatrix}\hskip 28.45274pt\text{v}_{3}=\begin{bmatrix}0&0&1\\ 0&0&0\\ 0&0&0\end{bmatrix}$
For $i\in\mathbb{N}$, consider the left invariant Riemannian metric $g_{i}$ on $H^{3}(\mathbb{R})$ for which
$\|\text{v}_{1}\|_{g_{i}}^{4}=\|\text{v}_{2}\|_{g_{i}}^{4}=\|\text{v}_{3}\|_{g_{i}}=1/i^{4}\hskip 56.9055pt\langle\text{v}_{j_{1}},\text{v}_{j_{2}}\rangle_{g_{i}}=0\text{ for }j_{1}\neq j_{2}\hskip 14.22636pt$
and let $X_{i}$ be the quotient of $(H^{3}(\mathbb{R}),g_{i})$ by the action of $H^{3}(\mathbb{Z})$. Then
${\text{diam}}(X_{i})\to 0\hskip 85.35826pt|\sec(X_{i})|\to 0$
Note that ${\text{rank}}(\pi_{1}(X_{i}))={\text{rank}}(H^{3}(\mathbb{Z}))=3$. If one wants to construct finite-sheeted covering spaces $X_{i}^{\prime}\to X_{i}$ with
${\text{diam}}(X_{i}^{\prime})\leq D\hskip 85.35826pt\mathcal{H}^{N}(X_{i}^{\prime})\geq\nu$
for some $D,\nu>0$, then one would need to unwind in all directions That is, if $u_{j}:=\exp(\text{v}_{j})\in H^{3}(\mathbb{Z})$ for $j\in\{1,2,3\}$ and $X_{i}^{\prime}=(H^{3}(\mathbb{R}),g_{i})/\Gamma_{i}$ for some subgroup $\Gamma_{i}\leq H^{3}(\mathbb{Z})$, then there is a sequence $N_{i}\in\mathbb{N}$ going to infinity at least as fast as $i$ for which
$u_{1}^{n_{1}}u_{2}^{n_{2}}u_{3}^{n_{3}}\notin\Gamma_{i}$
for all $n_{1},n_{2},n_{3}\in\mathbb{Z}$ with $|n_{1}|,|n_{2}|\leq N_{i}$, $|n_{3}|\leq N_{i}^{4}$. The projection $\rho:H^{3}(\mathbb{Z})\to\mathbb{Z}^{2}$ given by
$u_{1}\mapsto(1,0)\hskip 42.67912ptu_{2}\mapsto(0,1)\hskip 42.67912ptu_{3}\mapsto(0,0)$
sends $\Gamma_{i}$ to a lattice $\Lambda_{i}$ generated by a basis $\{\text{w}_{1,i},\text{w}_{2,i}\}$. If we denote by $A_{i}$ the area of the parallelogram spanned by $\text{w}_{1,i}$ and $\text{w}_{2,i}$, then $A_{i}\geq N_{i}^{4}$. This follows from the fact that if $g_{j,i}\in\Gamma_{i}\cap\rho^{-1}(\text{w}_{j,i})$, then $[g_{1,i},g_{2,i}]=u_{3}^{A_{i}}\in\Gamma_{i}$. From here, it is not hard to show that
${\text{diam}}(\mathbb{R}^{2}/\Lambda_{i})\geq\sqrt{A_{i}}/10\geq N_{i}^{2}/10.$
Finally, notice that the map $\overline{\rho}:H^{3}(\mathbb{R})\to\mathbb{R}^{2}$ given by
$\overline{\rho}\left(\begin{bmatrix}1&x&z\\ 0&1&y\\ 0&0&1\end{bmatrix}\right):=(x,y)$
induces an $i$-Lipschitz surjective map $X_{i}^{\prime}\to\mathbb{R}^{2}/\Lambda_{i}$. For $i$ large enough this implies
${\text{diam}}(X_{i}^{\prime})\geq{\text{diam}}(\mathbb{R}^{2}/\Lambda_{i})/i\geq N_{i}^{2}/10i\geq\sqrt{i}\to\infty$
which is a contradiction.
[…]
The missing piece to prove Conjecture 1.1 is Lemma 3. That is, to prove the conjecture it would be enough to show that if in addition to the conditions of such lemma, one has
  • $X$ is a non-collapsed $\text{RCD}(K,N)$ space.
  • $Y$ is an almost flat nilpotent Lie group of dimension $N$.
  • $f$ satisfies the estimate
    $(1-\delta)\,d(x,y)^{1+\delta}\leq d(f(x),f(y))\leq(1+\delta)\,d(x,y)\,\text{ for all }x,y\in B_{10}(p)$
  • $G_{Y}$ acts freely on $Y$.
then the produced $\tilde{f}$ can be chosen to be locally bi-Hölder (and not just a homotopy equivalence). Arguing as in [KM21, Theorem 3.3] one can construct a global homeomorphism $\tilde{f}:X\to Y$ that is locally bi-Hölder, but making this function equivariant seems to be a very interesting technical challenge.
In principle, one would need to choose the bump functions $\rho_{g}$ more carefully, but this may not be enough as one may need to also modify the function $f$ as it is done in the smooth setting by Huang in [H20]. Moreover, the exact same construction is possible as the groups we are dealing with are amenable, making the equivariant map constructed in [H20, Theorems 1.3, 1.8] the perfect candidate to be the desired map.
1.2. Collapsed $\text{RCD}(K,N)$ spaces
For an $\text{RCD}(K,N)$ space $(X,\mathsf{d},\mathfrak{m})$ with ${\text{diam}}(X)\leq\varepsilon(K,N)$, it follows from the proof of Theorem 1 that ${\text{rank}}(\pi_{1}(X))$ is at most the rectifiable dimension of $X$, which is an integer bounded above by $N$. One can ask whether the rigidity of Theorem 1 can be pushed to when ${\text{rank}}(\pi_{1}(X))$ equals the rectifiable dimension of $X$. This is not the case.
[…]
For each $n\geq 4$, $\varepsilon>0$, there is an $\text{RCD}(-1,n+2)$ space $(X,\mathsf{d},\mathfrak{m})$ of rectifiable dimension equal to $n$ with
${\text{diam}}(X)\leq\varepsilon\hskip 85.35826pt\pi_{1}(X)=\mathbb{Z}^{n}$
but no point of $X$ admits an open neighborhood homeomorphic to $\mathbb{R}^{n}$.
Proof.
This follows by performing the construction of [HNW23], which applies to general $4$-dimensional Riemannian manifolds, to the flat torus $\mathbb{T}^{4}$, and then multiplying by $\mathbb{T}^{n-4}$. ∎
It is however possible that if the dimension gap is small enough, the topological rigidity still holds.
[…]
For each $K\in\mathbb{R}$, $N\in\mathbb{N}$, there is $\varepsilon>0$ such that if $(X,\mathsf{d},\mathfrak{m})$ is an $\text{RCD}(K,N+\varepsilon)$ with ${\text{diam}}(X)\leq\varepsilon$ and ${\text{rank}}(\pi_{1}(X))=N$, then $X$ is homeomorphic to an infranilmanifold.
1.3. Fibrations
Theorem 1 is a step towards the following conjectures which would extend to $\text{RCD}(K,N)$ spaces work of Naber–Zhang [NZ16], Huang–Kong–Rong–Xu [HKRX20], and Huang–Wang [HW20].
[…]
Let $X$ be a geodesic space, $x\in X$, and $\varepsilon>0$. The fibered fundamental group $\Gamma_{\varepsilon}(x)$ is defined to be the image of the map $\pi_{1}(B_{\varepsilon}(x))\to\pi_{1}(B_{1}(x))$ induced by the inclusion.
[…]
There is $\varepsilon(K,N)>0$ such that the following holds. Let $X$ be a closed $m$-dimensional Riemannian manifold with
$|\sec(X)|\leq 1\hskip 85.35826pt{\text{inj}}(X)\geq 1$
and $(X_{i},\mathsf{d}_{i},\mathfrak{m}_{i})$ a sequence of $\text{RCD}(K,N)$ spaces converging in the Gromov–Hausdorff sense to $X$. Assume that for all $i$ large enough and any $x\in X_{i}$ one has
${\text{rank}}(\Gamma_{\varepsilon}(x))=N-m.$
Then for large enough $i$ there are continuous Gromov–Hausdorff approximations $X_{i}\to X$ that are locally trivial fibrations with infranilmanifold fiber.
[…]
Let $(X_{i},\mathsf{d}_{i},\mathfrak{m}_{i})$ be a sequence of compact $\text{RCD}(K,N)$ spaces that converges in the measured Gromov–Hausdorff sense to a closed $m$-dimensional Riemannian manifold $X$. Assume that for all $i$ one has
$\text{b}_{1}(X_{i})-\text{b}_{1}(X)=N-m.$
Then for large enough $i$ there are continuous Gromov–Hausdorff approximations $X_{i}\to X$ that are also locally trivial fibrations with fiber $\mathbb{T}^{N-m}$.
1.4. Overview
In Section 2 we cover the background material (or give corresponding references) and in Section 3 we prove the technical topological tools required for our main theorems.
In Section 4 we prove Theorem 1 by following the same strategy as Gromov in [G78] (see also [BK81] for a much more detailed explanation). First one needs to establish that if $(X,\mathsf{d},\mathfrak{m})$ is an $\text{RCD}(-\varepsilon,N)$ space with ${\text{diam}}(X)\leq\varepsilon$ and ${\text{rank}}(\pi_{1}(X))=N$, then the fundamental group of $X$ is the one of an infranilmanifold of dimension $N$. In particular, there is a finite index subgroup $\Gamma\leq\pi_{1}(X)$ which is isomorphic to a lattice in a simply connected nilpotent Lie group $Y$.
Furthermore, a large ball $B$ in the universal cover $\tilde{X}$ is both homeomorphic and Gromov–Hausdorff close to a ball in the Euclidean space $\mathbb{R}^{N}$, the same being true for a ball $B^{\prime}$ in $Y$ when equipped with an appropriate invariant Riemannian metric. Moreover, the homeomorphism $f:B\to B^{\prime}$ can be taken to be almost equivariant with respect to the local actions of the group $\Gamma$ on both balls.
Using the results from Section 3, via a center of mass construction, the map $f$ is extended to a global $\Gamma$-equivariant homotopy equivalence $\tilde{f}:\tilde{X}\to Y$. Since the Borel Conjecture holds for virtually nilpotent groups, this finishes the proof.
In Section 5 we fill the gap in the proof of Theorem 1.1 by adapting to the equivariant setting a topological stability result of Honda–Peng [HP23, Theorem 6.3].
Acknowledgement
The authors would like to thank Shouhei Honda, Raquel Perales, Xiaochun Rong and Daniele Semola for helpful discussions. Sergio Zamora holds a Postdoctoral Fellowship at Max Planck Institute for Mathematics at Bonn.
2. Preliminary material
2.1. Notation
For a metric space $X$, we denote its group of isometries by $\text{Iso}(X)$. For $p\in X$, $r>0$, we denote the open ball of radius $r$ around $p$ by $B_{r}^{X}(p)$ or $B_{r}(p)$, depending on whether we want to emphasize the metric space used.
For $N\geq 0$, we denote by $\mathcal{H}^{N}$ the $N$-dimensional Hausdorff measure.
For $N\in\mathbb{N}$, we denote by $\mathbb{T}^{N}$ the torus
$\{(x_{1},y_{1},\ldots,x_{N},y_{N})\in\mathbb{R}^{2N}\,|\,x_{j}^{2}+y_{j}^{2}=1\text{ for each }j\}$
equipped with its standard flat metric and the $N$-dimensional Hausdorff measure.
$H_{m}(\cdot)$ denotes singular homology with integer coefficients. Singular simplices, singular chains, and the boundary operator $\partial$ are as defined in [H02, Chapter 2].
[…]
2.2. $\text{RCD}(K,N)$ spaces
Throughout this text, we assume the reader is familiar with the basic theory of $\text{RCD}(K,N)$ spaces, including
  • Definition of Sobolev space $H^{1,2}$ and Laplace operator $\Delta$ on (infinitesimally Hilbertian) metric measure spaces.
  • Definition of $\text{RCD}(K,N)$ spaces, GH, pGH, and pmGH convergence.
  • Compactness of $\text{RCD}(K,N)$ spaces with respect to pmGH topology.
  • Definition of rectifiable dimension of an $\text{RCD}(K,N)$ space.
  • Existence of simply connected universal cover and $\text{RCD}(K,N)$ structure on it.
  • Closure of $\text{RCD}(K,N)$ spaces under quotients by compact and discrete groups of measure preserving isometries.
  • Equivariant Gromov–Hausdorff convergence of groups of isometries.
We refer the reader to [LV09, S06a, S06b, AGS14] for the first two items, and [EKS15] for an overview of different equivalent definitions of $\text{RCD}(K,N)$ spaces. We note that a large number of papers in the literature work with a condition known as $\text{RCD}^{\ast}(K,N)$, originally introduced in [AGMR15]. It is now known that this condition is equivalent to the $\text{RCD}(K,N)$ condition [CM21, L24].
The third item is covered for example in [LV09, GMS15]. The fourth item is covered in [BS23], where it is proven that the rectifiable dimension is well defined.
In [W24], it is shown that an $\text{RCD}(K,N)$ space $(X,\mathsf{d},\mathfrak{m})$ admits a simply-connected universal cover $\tilde{X}$, and in [MW19] an $\text{RCD}(K,N)$ structure on $\tilde{X}$ is constructed so that the covering map $\tilde{X}\to X$ is a local isomorphism of metric measure spaces.
In [GKMS18], it is shown that for an $\text{RCD}(K,N)$ space $(X,\mathsf{d},\mathfrak{m})$ and a group of measure preserving isometries $G\leq\text{Iso}(X)$, the metric quotient $X/G$ admits a natural $\text{RCD}(K,N)$ structure, provided $G$ is compact or discrete.
We refer to [G81, FY92] for the definition of equivariant Gromov–Hausdorff convergence (see also [BFT17] for a more recent exposition using ultralimits).
2.3. Volume convergence
Recall that an $\text{RCD}(K,N)$ space $(X,\mathsf{d},\mathfrak{m})$ is said to be non-collapsed if $N\in\mathbb{Z}$ and $\mathfrak{m}=\mathcal{H}^{N}$. In such a case, by the work of Mondino–Naber [MN19, Theorem 6.5], its rectifiable dimension is $N$. An important feature of non-collapsed $\text{RCD}(K,N)$ spaces is the continuity of volume, originally proven by Colding in the smooth setting [C97].
[…]
Let $(X_{i},\mathsf{d}_{i},\mathfrak{m}_{i},p_{i})$ be a sequence of pointed $\text{RCD}(K,N)$ spaces such that the sequence $(X_{i},p_{i})$ converges in the pointed Gromov–Hausdorff sense to a pointed non-collapsed $\text{RCD}(K,N)$ space $(X,p)$. Then
  1. (1)
    For $i$ large enough, $\mathfrak{m}_{i}=c_{i}\mathcal{H}^{N}$ for a sequence $c_{i}\in(0,\infty)$.
  2. (2)
    $\lim_{i\to\infty}\mathcal{H}^{N}(B_{1}^{X_{i}}(p_{i}))=\mathcal{H}^{N}(B_{1}^{X}(p)).$
Proof.
By the lower semi-continuity of the rectifiable dimension [K19, Theorem 1.5], the rectifiable dimension of $X_{i}$ is $N$ for $i$ large enough. Then the the first claim follows from [BGHZ23], and the second one from the work of De Philippis–Gigli [DPG18, Theorem 1.3]. ∎
[…]
Let $(X_{i},p_{i})$ be a sequence of proper geodesic spaces. We say a sequence of groups $W_{i}\leq\text{Iso}(X_{i})$ consists of small subgroups if for each $R>0$ we have
$\lim_{i\to\infty}\,\sup_{g\in W_{i}}\,\sup_{x\in B_{R}(p_{i})}d(gx,x)=0.$
Equivalently, the groups $W_{i}$ are small if they converge to the trivial group in the equivariant Gromov–Hausdorff sense.
[…]
Let $(X_{i},\mathsf{d}_{i},\mathfrak{m}_{i},p_{i})$ be a sequence of pointed $\text{RCD}(K,N)$ spaces such that the sequence $(X_{i},p_{i})$ converges in the pointed Gromov–Hausdorff sense to a pointed non-collapsed $\text{RCD}(K,N)$ space $(X,p)$. Then any sequence of small subgroups $W_{i}\leq\text{Iso}(X_{i})$ is eventually trivial.
Proof.
Working by contradiction, after passing to a subsequence, we assume $W_{i}$ is non-trivial for each $i$. For $i$ large enough, the closure $\overline{W_{i}}$ is a non-trivial compact Lie group, so it contains a non-trivial finite group $F_{i}\leq\overline{W_{i}}$. By Theorem 2.3, the groups $\text{Iso}(X_{i})$ are measure preserving, so the space $X_{i}/F_{i}$ admits a measure that makes it an $\text{RCD}(K,N)$ space, and since the groups $F_{i}$ are small, the sequence $(X_{i}/F_{i},[p_{i}])$ also converges in the pointed Gromov–Hausdorff sense to $(X,p)$. However, applying Theorem 2.3(2) twice we get
$\mathcal{H}^{N}(B_{1}^{X}(p))=\lim_{i\to\infty}\mathcal{H}^{N}(B_{1}^{X_{i}/F_{i}}([p_{i}]))=\lim_{i\to\infty}\frac{1}{|F_{i}|}\mathcal{H}^{N}(B_{1}^{X_{i}}(p_{i}))\leq\frac{1}{2}\mathcal{H}^{N}(B_{1}^{X}(p)),$
which is a contradiction. ∎
2.4. Topological stability
The following is a special case of [KM21, Theorem 2.9].
[…]
Let $(X_{i},\mathsf{d}_{i},\mathfrak{m}_{i},p_{i})$ be a sequence of pointed $\text{RCD}(-\frac{1}{i},N)$ spaces such that the sequence $(X_{i},p_{i})$ converges in the pointed Gromov–Hausdorff sense to $(\mathbb{R}^{N},0)$. Then there is a sequence of pointed Gromov–Hausdorff approximations $f_{i}:(X_{i},p_{i})\to(\mathbb{R}^{N},0)$ such that for all $R>\delta>0$, the restriction of $f_{i}$ to $B_{R}(p_{i})$ is a homeomorphism with its image for $i$ large enough depending on $R,\delta$, and
$B_{R-\delta}(0)\subset f_{i}(B_{R}(p_{i})).$
[…]
Let $(X_{i},\mathsf{d}_{i},\mathfrak{m}_{i},p_{i})$ be a sequence of pointed $\text{RCD}(-\frac{1}{i},N)$ spaces such that the sequence $(X_{i},p_{i})$ converges in the pointed Gromov–Hausdorff sense to $(\mathbb{R}^{N},0)$. If there are groups $\Gamma_{i}\leq\text{Iso}(X_{i})$ with
$\limsup_{i\to\infty}\,{\text{diam}}(X_{i}/\Gamma_{i})<\infty,$
then $X_{i}$ is a topological manifold of dimension $N$ for $i$ large enough.
Proof.
Let $R>\limsup_{i}{\text{diam}}(X_{i}/\Gamma_{i})$. By Theorem 2.4, for $i$ large enough each point in $B_{R}(p_{i})$ has a neighborhood homeomorphic to $\mathbb{R}^{N}$. Since also for $i$ large enough and each $x\in X_{i}$ there is an isometry $g\in\Gamma_{i}$ with $gx\in B_{R}(p_{i})$ the result follows. ∎
[…]
Let $(X_{i},\mathsf{d}_{i},\mathfrak{m}_{i},p_{i})$ be a sequence of pointed $\text{RCD}(-\frac{1}{i},N)$ spaces such that the sequence $(X_{i},p_{i})$ converges in the pointed Gromov–Hausdorff sense to $(\mathbb{R}^{N},0)$. Then for any $\varepsilon$, the following holds for $i$ large enough:
  • For all $r\in(0,1]$, $x\in B_{1}(p_{i})$, the ball $B_{r}(x)$ is contractible in $B_{r+\varepsilon}(x)$.
Proof.
Let $f_{i}:X_{i}\to X$ be the Gromov–Hausdorff approximations given by Theorem 2.4. Assuming the result fails, after passing to a subsequence, there are sequences $x_{i}\in B_{1}(p_{i})$, $r_{i}\in(0,1]$ for which the ball $B_{r_{i}}(x_{i})$ is not contractible in $B_{r_{i}+\varepsilon}(x_{i})$. After passing further to a subsequence, we can assume $r_{i}\to r$ for some $r\in[0,1]$ and $f_{i}(x_{i})\to x$ for some $x\in B_{1}(0)$.
Then for $i$ large enough, $B_{r+\varepsilon/4}(x_{i})$ is not contractible in $B_{r+3\varepsilon/4}(x_{i})$. However, this is false, as for $i$ large enough one has
$f_{i}(B_{r+\varepsilon/4}(x_{i}))\subset B_{r+\varepsilon/2}(x)\subset f_{i}(B_{r+3\varepsilon/4}(x_{i})),$
so we can pull back via $f_{i}$ a nullhomotopy of $B_{r+\varepsilon/2}(x)$. ∎
2.5. Analysis and Spectral Estimates
Let $(X,\mathsf{d},\mathfrak{m})$ be a compact $\text{RCD}(K,N)$ space of diameter $\leq D$. The spectrum of its Laplacian $\Delta$ is discrete and non-positive, and one has an orthogonal direct sum decomposition
$H^{1,2}(X)=\bigoplus_{\lambda\geq 0}\{f\in H^{1,2}(X)|\Delta f=-\lambda f\}$
For an eigenfunction $f\in H^{1,2}(X)$ with $\Delta f=-\lambda f$, one has
$\|f\|_{H^{1,2}}^{2}=(1+\lambda)\|f\|_{L^{2}}^{2}$
Hence by [AHPT21, Proposition 7.1] and the discussion preceding it, there is $C(K,N,D)>0$ for which
(2.1) $\hskip 17.07182pt\|f\|_{L^{\infty}}\leq C\lambda^{(N-2)/4}\|f\|_{H^{1,2}}\hskip 34.14322pt\|\nabla f\|_{L^{\infty}}\leq C\lambda^{N/4}\|f\|_{H^{1,2}}$
The Laplacian spectrum is continuous with respect to measured Gromov–Hausdorff convergence. This continuity was originally conjectured for Riemannian manifolds by Fukaya [F87], proven in that context and more general ones by Cheeger–Colding [CC00], and extended to $\text{RCD}(K,N)$ spaces by Gigli–Mondino–Savare [GMS15] (see also [AH18] for an extension in non-compact spaces).
[…]
In a metric space $Z$, we say a sequence of boundedly finite measures $\mathfrak{m}_{i}$ converges weakly to a measure $\mathfrak{m}$ if for any bounded continuous function $f:Z\to\mathbb{R}$ with bounded support, one has
$\int_{Z}f\,\text{d}\mathfrak{m}_{i}\to\int_{Z}f\,\text{d}\mathfrak{m}$
[…]
Let $(X_{i},\mathsf{d}_{i},\mathfrak{m}_{i})$ be a sequence of $\text{RCD}(K,N)$ spaces converging in the measured Gromov–Hausdorff sense to an $\text{RCD}(K,N)$ space $(X,\mathsf{d},\mathfrak{m})$. This means there is a complete metric space $Z$ and isometric embeddings $\iota_{i}:X_{i}\to Z$, $\iota:X\to Z$, such that the sequence $\iota_{i}(X_{i})$ converges in the Hausdorff topology to $\iota(X)$ and the measures $(\iota_{i})_{\ast}\mathfrak{m}_{i}$ converge weakly to $(\iota)_{\ast}\mathfrak{m}$. Given a sequence $f_{i}\in H^{1,2}(X_{i})$, we say it converges in $H^{1,2}$ to $f\in H^{1,2}(X)$ if the sequence $(\iota_{i})_{\ast}(f_{i}\mathfrak{m}_{i})$ converges weakly to $(\iota)_{\ast}(f\mathfrak{m})$ and
$\lim_{i\to\infty}\|f_{i}\|_{H^{1,2}}=\|f\|_{H^{1,2}}$
We say $f_{i}$ converges uniformly to $f$ if for any choice of points $x_{i}\in X_{i}$, $x\in X$ with $\iota_{i}(x_{i})\to\iota(x)$, one has $f_{i}(x_{i})\to f(x)$.
[…]
Let $(X_{i},\mathsf{d}_{i},\mathfrak{m}_{i})$ be a sequence of compact $\text{RCD}(K,N)$ spaces that converges in the measured Gromov–Hausdorff sense to a compact $\text{RCD}(K,N)$ space $(X,\mathsf{d},\mathfrak{m})$. Then
  1. (1)
    For each sequence of eigenfunctions $f_{i}\in H^{1,2}(X_{i})$ with
    $\Delta f_{i}=-\lambda_{i}f_{i}\hskip 28.45274pt\limsup_{i\to\infty}\lambda_{i}<\infty\hskip 28.45274pt\limsup_{i\to\infty}\|f_{i}\|_{H^{1,2}}<\infty$
    there is an eigenfunction $f\in H^{1,2}(X)$ with $\Delta f=-\lambda f$ such that after passing to a subsequence, $\lambda_{i}\to\lambda$ and $f_{i}$ converges to $f$ both uniformly and in $H^{1,2}$.
  2. (2)
    For each eigenfunction $f\in H^{1,2}(X)$ with $\Delta f=-\lambda f$, there is a sequence of eigenfunctions $f_{i}\in H^{1,2}(X_{i})$ with $\Delta f_{i}=-\lambda_{i}f_{i}$ such that $\lambda_{i}\to\lambda$ and $f_{i}$ converges to $f$ both uniformly and in $H^{1,2}$.
[…]
Let $(X_{i},\mathsf{d}_{i},\mathfrak{m}_{i})$ be a sequence of compact $\text{RCD}(K,N)$ spaces that converges in the measured Gromov–Hausdorff sense to a compact $\text{RCD}(K,N)$ space $(X,\mathsf{d},\mathfrak{m})$. Then the sequence of Laplace spectra of the spaces $X_{i}$ counted with multiplicities as measures on $\mathbb{R}$ converge weakly to the Laplace spectrum of $X$.
2.6. Canonical homeomorphisms
Eigenfunctions provide means of constructing bi-Hölder homeomorphisms between suitable non-collapsed $\text{RCD}(K,N)$ spaces [HP23, Theorem 1.2.2] (see also [WZ23]).
[…]
Let $(X_{i},\mathsf{d}_{i},\mathfrak{m}_{i})$ be a sequence of compact $\text{RCD}(K,N)$ spaces. A sequence of maps $\Phi_{i}:X_{i}\to\mathbb{R}^{k}$ is said to be equi-regular if for any smooth function $f\in C^{\infty}(\mathbb{R}^{k})$, the composition $f\circ\Phi_{i}:X_{i}\to\mathbb{R}$ is in the domain of the Laplacian for all $i$ and
$\sup_{i}\|\Delta(f\circ\Phi_{i})\|_{\infty}<\infty$
[…]
Let $(X_{i},\mathsf{d}_{i},\mathfrak{m}_{i})$ be a sequence of compact $\text{RCD}(K,N)$ spaces that converges in the measured Gromov–Hausdorff sense to a compact $\text{RCD}(K,N)$ space $(X,\mathsf{d},\mathfrak{m})$. Assume $(X,\mathsf{d})$ is isometric to a smooth $N$-dimensional Riemannian manifold, and consider a smooth isometric embedding $\Phi:X\to\mathbb{R}^{k}$, together with an open neighborhood $U\subset\mathbb{R}^{k}$ and a $C^{1,1}$-Riemannian submersion $\pi:U\to\Phi(X)$ with $\pi|_{\Phi(X)}=\text{Id}_{\Phi(X)}$. Then if $\Phi_{i}=(\varphi_{1,i},\ldots,\varphi_{k,i}):X_{i}\to\mathbb{R}^{k}$ is a sequence of equi-regular maps converging uniformly to $\Phi$ such that the family of functions $\Delta\varphi_{j,i}$ is equi-Lipschitz, then the sequence of maps
$F_{i}:=(\Phi^{-1}\circ\pi\circ\Phi_{i}):X_{i}\to X$
satisfy
$(1-\varepsilon_{i})\,\mathsf{d}_{i}(x,y)^{1+\varepsilon_{i}}\leq\mathsf{d}(F_{i}(x),F_{i}(y))\leq(1+\varepsilon_{i})\,\mathsf{d}_{i}(x,y)\,\text{ for all }x,y\in X_{i}$
for some $\varepsilon_{i}\to 0$. Furthermore, if the spaces $X_{i}$ are smooth Riemannian manifolds, and $\Phi_{i}$ is smooth for each $i$, then $F_{i}$ is a diffeomorphism for $i$ large enough.
[…]
Let $(X_{i},\mathsf{d}_{i},\mathfrak{m}_{i})$ be a sequence of compact $\text{RCD}(K,N)$ spaces of uniformly bounded diameter and $\varphi_{j,i}\in H^{1,2}(X_{i})$ a family of eigenfunctions with $\Delta\varphi_{j,i}=-\lambda_{j,i}\varphi_{j,i}$ for $j\in\{1,\ldots,k\}$ and assume
$\sup_{i,j}\,\lambda_{j,i}<\infty\hskip 85.35826pt\sup_{i,j}\,\|\varphi_{j,i}\|_{2}<\infty$
Consider the map $\Phi_{i}=(\varphi_{1,i},\ldots,\varphi_{k,i}):X_{i}\to\mathbb{R}^{k}$. Since for any smooth function $f\in C^{\infty}(\mathbb{R}^{k})$, one has
$\Delta(f\circ\Phi_{i})=\sum_{j=1}^{k}\left[\frac{\partial f}{\partial x_{j}}\circ\Phi_{i}\right]\Delta\varphi_{j,i}+\sum_{j,\ell=1}^{k}\left[\frac{\partial^{2}f}{\partial x_{j}\partial x_{\ell}}\circ\Phi_{i}\right]\langle\nabla\varphi_{j,i},\nabla\varphi_{\ell,i}\rangle,$
then by (2.1), the maps $\Phi_{i}=(\varphi_{1,i},\ldots,\varphi_{k,i}):X_{i}\to\mathbb{R}^{k}$ are equi-regular and equi-Lipschitz.
2.7. Group theory
Recall that a finitely generated group $\Lambda$ is said to be polycyclic if there is a sequence of subgroups
$1=\Lambda_{0}\trianglelefteq\cdots\trianglelefteq\Lambda_{m}=\Lambda$
with $\Lambda_{j}/\Lambda_{j-1}$ cyclic for each $j$. In such a case, ${\text{rank}}(\Lambda)$ is defined to be the number of $j$’s for which the quotient $\Lambda_{j}/\Lambda_{j-1}$ is isomorphic to $\mathbb{Z}$. By Schreier Theorem [F67] this number is well defined. The following properties come directly from the definition.
  1. R.1
    If $\Lambda^{\prime}\leq\Lambda$ is a finite index subgroup then ${\text{rank}}(\Lambda^{\prime})={\text{rank}}(\Lambda)$.
  2. R.2
    For a short exact sequence
    $1\to K\to\Lambda\to\Gamma\to 1$
    with $\Lambda$ a polycyclic group, we have
    ${\text{rank}}(\Lambda)={\text{rank}}(K)+{\text{rank}}(\Gamma).$
For an arbitrary finitely generated group $G$, ${\text{rank}}(G)$ is defined as the infimum of ${\text{rank}}(\Lambda)$ among finite index polycyclic subgroups $\Lambda\leq G$. By R.1 if $G$ is polycyclic there is no conflict between the distinct definitions of ${\text{rank}}(G)$.
[…]
Let $G$ be a group, $\Gamma$ a polycyclic group, and $\theta:G\to\Gamma$ a surjective morphism with finite kernel. Then ${\text{rank}}(G)={\text{rank}}(\Gamma)$.
Proof.
Let $\Lambda\leq G$ be the kernel of the adjoint action on ${\text{Ker}}(\theta)\trianglelefteq G$. Then there is a short exact sequence
$1\to{\text{Ker}}(\theta)\cap\Lambda\to\Lambda\to\theta(\Lambda)\to 1$
${\text{Ker}}(\theta)\cap\Lambda$ is finite abelian, hence polycyclic. Since ${\text{Ker}}(\theta)\cap\Lambda$ and $\theta(\Lambda)$ are polycyclic, so is $\Lambda$. Then by R.1 and R.2;
${\text{rank}}(G)={\text{rank}}(\Lambda)={\text{rank}}(\theta(\Lambda))={\text{rank}}(\Gamma).$
The ensuing lemma follows from classic simultaneous diagonalization.
[…]
Let $G$ be an abelian group acting by linear isometries on a finite dimensional real vector space $V$ of dimension $2N$. Assume $V=\oplus_{\ell=1}^{m}E_{\ell}$ for some $G$-invariant subspaces $E_{\ell}$. Then $V$ admits a (not necessarily unique!) decomposition $V=\oplus_{j=1}^{N}W_{j}$ with each $W_{j}$ 2-dimensional, $G$-invariant, and satisfying
$W_{j}=\bigoplus_{\ell=1}^{m}\,(E_{\ell}\cap W_{j})$
[…]
For a group $G$ and two subgroups $H_{1},H_{2}\leq G$, we denote by $[H_{1},H_{2}]\leq G$ the subgroup generated by elements of the form $h_{1}h_{2}h_{1}^{-1}h_{2}^{-1}$ with $h_{1}\in H_{1}$, $h_{2}\in H_{2}$. We set $G^{(0)}:=G$, and define inductively $G^{(j+1)}:=[G^{(j)},G]$. $G$ is said to be nilpotent if $G^{(s)}$ is the trivial group for some $s\in\mathbb{N}$.
2.8. Limits of almost homogeneous spaces
In a portion of this paper, we follow the construction performed in [Z24], which is an adaptation to metric spaces of the construction by Breuillard–Green–Tao in [BGT12]. The following lemma establishes that discrete groups that act almost transitively on geodesic spaces are approximately virtually nilpotent.
[…]
Let $(X_{i},p_{i})$ be a sequence of pointed proper geodesic spaces that converges to a pointed proper semi-locally-simply-connected geodesic space $(X,p)$ in the pointed Gromov–Hausdorff sense, and $G_{i}\leq\text{Iso}(X_{i})$ a sequence of discrete groups with ${\text{diam}}(X_{i}/G_{i})\to 0$. Then there exists $s\in\mathbb{N}$ and a sequence of finite-index subgroups $G_{i}^{\prime}\leq G_{i}$ with
(2.2) $\lim_{i\to\infty}\,\,\,\sup_{x\in X_{i}}\,\sup_{g\in(G_{i}^{\prime})^{(s)}}d(gx,x)=0$
(2.3) $\limsup_{i\to\infty}\,[G_{i}:G_{i}^{\prime}]<\infty$
Proof.
By [Z24, Theorem 4.1], $X$ is a nilpotent Lie group equipped with a sub-Finsler metric. Sub-Finsler metrics are locally doubling [NSW85], so [Z24, Remark 3.4] applies and [Z24, Lemma 3.2] provides the desired groups $G_{i}^{\prime}$. ∎
It turns out the groups $G_{i}^{\prime}$ act in some sense almost by translations. This is encoded in the following lemma.
[…]
Let $(X_{i},p_{i})$, $(X,p)$, $G_{i}$, $G_{i}^{\prime}$, be as in Lemma 2.8. Then after passing to a subsequence the groups $G_{i}^{\prime}$ converge in the equivariant Gromov–Hausdorff sense to a connected nilpotent group $\Gamma\leq\text{Iso}(X)$ acting freely and transitively.
Proof.
By precompactness [FY92, Proposition 3.6], after passing to a subsequence the groups $G_{i}^{\prime}$ converge to a group $\Gamma\leq\text{Iso}(X)$. By [Z24, Theorem 4.1] and its proof, $\Gamma$ is a nilpotent Lie group acting transitively. By [Z24, Proposition 5.2] and its proof, $\Gamma$ is connected and acts freely. ∎
[…]
Let $(X_{i},p_{i})$, $(X,p)$, $G_{i}$, $G_{i}^{\prime}$, be as in Lemma 2.8. If the groups $G_{i}$ are such that any sequence of small subgroups is eventually trivial, then $G_{i}^{\prime}$ acts freely for $i$ large enough.
Proof.
Assuming the contrary, after passing to a subsequence there are $x_{i}\in X_{i}$ with non-trivial stabilizers $K_{i}\leq G_{i}^{\prime}$. Notice that since the groups $G_{i}^{\prime}$ act almost transitively, we can assume $x_{i}\in B_{1}(p_{i})$ for $i$ large enough. After again passing to a subsequence, $x_{i}$ converges to a point $x\in X$ and the groups $K_{i}$ converge to a group $K\leq\Gamma$ contained in the stabilizer of $x$. By Lemma 2.8, $K$ is trivial, so the groups $K_{i}$ are small, contradicting the hypothesis. ∎
Lemma 2.8 is particularly powerful when paired with the following classic tool. It follows from the proof of [T02, Corollary 4.2.7].
[…]
Let $\Gamma\leq\text{Iso}(\mathbb{R}^{n})$ be a connected nilpotent group of isometries acting transitively. Then $\Gamma$ is the group of translations in $\mathbb{R}^{n}$.
Lemma 2.8 below gives an explicit description of the groups $G_{i}^{\prime}$.
[…]
Let $G$ be a group, $u_{1},u_{2},\ldots,u_{r}\in G$, and $N_{1},N_{2},\ldots,N_{r}$ $\in\mathbb{R}^{+}$. The set $P(u_{1},\ldots,u_{r};N_{1},\ldots,N_{r})\subset G$ is defined to be the set of elements that can be expressed as words in the $u_{i}$’s and their inverses such that the number of appearances of $u_{i}$ and $u_{i}^{-1}$ is not more than $N_{i}$. We then say that $P(u_{1},\ldots,u_{r};N_{1},\ldots,N_{r})$ is a nilprogression in $C$-normal form for some $C>0$ if it also satisfies the following properties:
  1. N.1
    For all $1\leq i\leq j\leq r$, and all choices of signs, we have
    $[u_{i}^{\pm 1},u_{j}^{\pm 1}]\in P\left(u_{j+1},\ldots,u_{r};\dfrac{CN_{j+1}}{N_{i}N_{j}},\ldots,\dfrac{CN_{r}}{N_{i}N_{j}}\right).$
  2. N.2
    The expressions $u_{1}^{n_{1}}\ldots u_{r}^{n_{r}}$ represent distinct elements as $n_{1},\ldots,n_{r}$ range over the integers with $|n_{1}|\leq N_{1}/C,\ldots,|n_{r}|\leq N_{r}/C$.
  3. N.3
    One has
    $\frac{1}{C}(2\lfloor N_{1}\rfloor+1)\cdots(2\lfloor N_{r}\rfloor+1)\leq|P|\leq C(2\lfloor N_{1}\rfloor+1)\cdots(2\lfloor N_{r}\rfloor+1).$
For a nilprogression $P$ in $C$-normal form, and $\varepsilon\in(0,1)$, the set $P(u_{1},\ldots,u_{r};$ $\varepsilon N_{1},\ldots,\varepsilon N_{r})$ also satisfies N.1 and N.2, and we denote it by $\varepsilon P$. We define the thickness of $P$ as the minimum of $N_{1},\ldots,N_{r}$ and we denote it by $\text{thick}(P)$. The set $\{u_{1}^{n_{1}}\ldots u_{r}^{n_{r}}||n_{i}|\leq N_{i}/C\}$ is called the grid part of $P$, and is denoted by $G(P)$.
[…]
Let $(X_{i},p_{i})$, $(X,p)$, $G_{i}$, $G_{i}^{\prime}$ be as in Lemma 2.8 and $n$ the topological dimension of $X$. Then for $i$ large enough, there are torsion-free nilpotent groups $\tilde{\Gamma}_{i}$ generated by elements $\tilde{u}_{1,i},\ldots,\tilde{u}_{n,i}\in\tilde{\Gamma}_{i}$, and $N_{1,i},\ldots,N_{n,i}\in\mathbb{R}^{+}$ with the following properties:
  1. (1)
    The set map $\alpha_{i}:\mathbb{Z}^{n}\to\tilde{\Gamma}_{i}$ given by
    $\alpha_{i}(m_{1},\ldots,m_{n})=\tilde{u}_{1,i}^{m_{1}}\cdots\tilde{u}_{n,i}^{m_{n}}$
    is a bijection. Moreover, there are polynomials $Q_{i}:\mathbb{R}^{n}\times\mathbb{R}^{n}\to\mathbb{R}^{n}$ of degree $\leq d(n)$ such that the product in $\tilde{\Gamma}_{i}$ is given by
    $g_{1}\cdot g_{2}=\alpha_{i}(Q_{i}(\alpha_{i}^{-1}(g_{1}),\alpha_{i}^{-1}(g_{2}))).$
  2. (2)
    There are small normal subgroups $W_{i}\trianglelefteq G_{i}^{\prime}$ and surjective group morphisms
    $\Phi_{i}:\tilde{\Gamma}_{i}\to\Gamma_{i}:=G_{i}^{\prime}/W_{i}$
    with ${\text{Ker}}(\Phi_{i})$ containing an isomorphic copy of $\pi_{1}(X)$.
  3. (3)
    There is $C>0$ such that if $u_{j,i}:=\Phi_{i}(\tilde{u}_{j,i})$ for each $j\in\{1,\ldots,n\}$, the set
    $P_{i}:=P(u_{1,i},\ldots,u_{n,i};N_{1,i},\ldots,N_{n,i})\subset\Gamma_{i}$
    is a nilprogression in $C$-normal form with $\text{thick}(P_{i})\to\infty$.
  4. (4)
    For each $\varepsilon>0$ there is $\delta>0$ such that
    $G(\delta P_{i})\subset\{g\in\Gamma_{i}\,|\,$ $d(g[p_{i}],[p_{i}])\leq\varepsilon\},$
    $\{g\in\Gamma_{i}\,|\,d(g[p_{i}],[p_{i}$ $])\leq\delta\}\subset G(\varepsilon P_{i})$
    for $i$ large enough, where we are considering the action of $\Gamma_{i}$ on $X_{i}/W_{i}$,
Proof.
In [Z24, Section 6] first the small subgroups $W_{i}\trianglelefteq G_{i}^{\prime}$ are constructed (using the escape norm from [BGT12]) in such a way that the quotients $G_{i}^{\prime}/W_{i}$ have no small subgroups. Then in [BGT12, Section 9] the nilprogressions $P_{i}\subset\Gamma_{i}$ are obtained via a short basis construction. (3) and the first item of (4) follow from [BGT12, Theorem 9.3] and its proof, while the second item of (4) is proven in [Z24, Theorem 7.1]. In [Z24, Section 8], (1) is obtained by an application of Malcev Theorem, and (2) follows from [Z24, Proposition 9.1]. Note that although the results in [BGT12, Z24] are stated in terms of ultralimits and non-standard analysis, they are equivalent to the above formulations by Łos’s Theorem (see for example [BGT12, Appendix C]). ∎
[…]
Combining Lemma 2.8 with Lemma 2.7 and R.2 one obtains
(2.4) $\begin{split}{\text{rank}}(G_{i}^{\prime})&={\text{rank}}(\Gamma_{i})\\ &={\text{rank}}(\tilde{\Gamma}_{i})-{\text{rank}}({\text{Ker}}(\Phi_{i}))\\ &\leq n.\end{split}$
If equality holds throughout (2.4), then ${\text{rank}}({\text{Ker}}(\Phi_{i}))=0$, and since the groups $\tilde{\Gamma}_{i}$ are torsion-free, that would imply ${\text{Ker}}(\Phi_{i})=0$, so $X$ is simply connected and $\tilde{\Gamma}_{i}=\Gamma_{i}$.
[…]
If the groups $G_{i}^{\prime}$ admit no small subgroups, then $\Gamma_{i}=G_{i}^{\prime}$ for $i$ large enough.
3. Topological Lemmas
In general, when two pointed proper metric spaces $(X,p)$, $(Y,q)$ are pointed Gromov–Hausdorff close, there is no global approximation $X\to Y$. However, when a group $G$ acts almost transitively on both spaces and the triples $(X,p,G)$, $(Y,q,G)$ are equivariantly Gromov–Hausdorff close, it is possible to construct a global almost $G$-equivariant approximation $X\to Y$.
[…]
For each $\varepsilon>0$, there is $\delta>0$ such that the following holds. Let $X$, $Y$ be proper geodesic spaces, $G_{X}\leq\text{Iso}(X)$, $G_{Y}\leq\text{Iso}(Y)$ discrete groups, $p\in X$, $q\in Y$, a group isomorphism $\varphi:G_{X}\to G_{Y}$, and a map $f:B_{10}(p)\to B_{10}(q)$ such that $f(p)=q$ and
  1. (a)
    ${\text{diam}}(X/G_{X}),$ ${\text{diam}}(Y/G_{Y})\leq\delta$.
  2. (b)
    For all $x_{1},x_{2}\in B_{10}(p)$,
    $|d(f(x_{1}),f(x_{2}))-d(x_{1},x_{2})|\leq\delta.$
  3. (c)
    If $x\in B_{10}(p)$, $h\in G_{Y}$ are such that $hf(x)\in B_{5}(q)$, then
    $\varphi^{-1}(h)x\in B_{10}(p).$
  4. (d)
    If $x\in B_{10}(p)$, $g\in G_{X}$ are such that $gx\in B_{10}(p)$, then
    $d(f(gx),\varphi(g)f(x))\leq\delta.$
Then there is $\tilde{f}:X\to Y$ such that
  1. (1)
    For all $x\in X$, $g\in G_{X}$,
    $d(\phi(g)\tilde{f}(x),\tilde{f}(gx))\leq\varepsilon.$
  2. (2)
    For all $x_{1},x_{2}\in X$ with $\min\{d(x_{1},x_{2}),d(\tilde{f}(x_{1}),\tilde{f}(x_{2}))\}\leq 2$ one has
    $|d(\tilde{f}(x_{1})\tilde{f}(x_{2}))-d(x_{1},x_{2})|\leq\varepsilon.$
  3. (3)
    For all $y\in Y$, there is $x\in X$ with $d(\tilde{f}x,y)\leq\varepsilon$.
Moreover, if
(3.1) $Y\text{ is Riemannian,}\,\,\,\,|\sec(Y)|\leq 1,\,\,\,\,{\text{inj}}(Y)\geq 1,$
and $f$ is continuous, then $\tilde{f}$ can be taken to be continuous and equivariant; that is, for all $x\in X$, $g\in G_{X}$,
$\varphi(g)\tilde{f}(x)=\tilde{f}(gx).$
Proof.
For each $g\in G_{X}$, set $\theta_{g}:X\to\mathbb{R}$ as
$\theta_{g}(x):=\max\{0,1-d(x,gp)\}.$
Then define $\Theta:X\to\mathbb{R}$ as
$\Theta(x)=\sum_{g\in G_{X}}\theta_{g}(x),$
which by (a) is always positive. Also for each $g\in G_{X}$ define $\rho_{g}:X\to\mathbb{R}$ as
$\rho_{g}(x):=\theta_{g}(x)/\Theta(x),$
and with this, the map $\mu:X\to\text{Prob}(Y)$ as
$\mu_{x}:=\sum_{g\in G_{X}}\rho_{g}(x)[\varphi(g)(f(g^{-1}x))],$
where for each $y\in Y$, we denote by $[y]\in\text{Prob}(Y)$ the delta measure supported at $y$. Notice that $\rho_{g}(x)\neq 0$ means that $g^{-1}x\in B_{1}(p)$, so $\mu_{x}$ is well defined even if $f(g^{-1}x)$ is not for most $g\in G_{X}$. It follows from the definitions that
$\rho_{g}(hx)=\rho_{h^{-1}g}(x)\text{ for all }x\in X,\,g,h\in G_{X}.$
Then
(3.2) $\begin{split}\mu_{hx}&=\sum_{g\in G_{X}}\rho_{g}(hx)[\varphi(g)(f(g^{-1}hx))]\\ &=\sum_{g\in G_{X}}\rho_{h^{-1}g}(x)[\varphi(g)(f((h^{-1}g)^{-1}x))]\\ &=\sum_{u\in G_{X}}\rho_{u}(x)[\varphi(hu)(f(u^{-1}x))]\\ &=\varphi(h)_{\ast}\sum_{u\in G_{X}}\rho_{u}(x)[\varphi(u)(f(u^{-1}x))]\\ &=\varphi(h)_{\ast}\mu_{x},\end{split}$
so $\mu$ is an equivariant map. Also, if $x\in X$, $g,h\in G_{X}$ are such that $\rho_{g}(x),\rho_{h}(x)>0$, then $g^{-1}(x),h^{-1}(x)\in B_{1}(p)$, and by condition (d) we have
$d(\varphi(g)f($ $g^{-1}x),\varphi(h)f(h^{-1}x))=d(f(g^{-1}x),\varphi(g^{-1}h)f(h^{-1}x))$
$=d(f((g^{-1}h)h^{-1}x),\varphi(g^{-1}h)f(h^{-1}x))$
$\leq\delta,$
so
(3.3) ${\text{diam}}({\text{supp}}(\mu_{x}))\leq\delta\text{ for all }x\in X.$
Then define $\tilde{f}:X\to Y$ as any map with
(3.4) $\tilde{f}(x)\in{\text{supp}}(\mu_{x}).$
By (3.3), different choices for $\tilde{f}$ yield functions that are uniformly $\delta$-close, and (1) follows from (3.2). Now if $x_{1},x_{2}\in X$ are such that $d(x_{1},x_{2})\leq 5,$ take $g_{1},g_{2}\in G_{X}$ with $g^{-1}_{1}x_{1},g^{-1}_{2}x_{2}\in B_{1}(p)$. Then $g_{1}^{-1}x_{2}\in B_{10}(p)$ so by (b) and (d) we have
$d(\varphi(g_{1})f($ $g_{1}^{-1}x_{1}),\varphi(g_{2})f(g_{2}^{-1}x_{2}))=d(f(g_{1}^{-1}x_{1}),\varphi(g_{1}^{-1}g_{2})f(g_{2}^{-1}x_{2}))$
$=d(f(g_{1}^{-1}x_{1}),f(g_{1}^{-1}x_{2}))\pm\delta$
$=d(g_{1}^{-1}x_{1},g_{1}^{-1}x_{2})\pm 2\delta$
$=d(x_{1},x_{2})\pm 2\delta.$
Combining this with (3.3) we obtain
(3.5) $|d(\tilde{f}(x_{1}),\tilde{f}(x_{2}))-d(x_{1},x_{2})|\leq 4\delta.$
On the other hand, if $x_{1},x_{2}$ are such that $d(\tilde{f}(x_{1}),\tilde{f}(x_{2}))\leq 2$, then there are $g_{1},g_{2}\in G_{X}$ such that $\rho_{g_{1}}(x_{1}),\rho_{g_{2}}(x_{2})>0$, and
$d(\varphi(g_{1})f(g_{1}^{-1}x_{1}),\varphi(g_{2})f(g_{2}^{-1}x_{2}))\leq 3.$
Using (b) and the fact that $f(p)=q$, this implies
$d(\varphi(g_{2}^{-1}g_{1})f(g_{1}^{-1}x_{1}),q)\leq 5,$
so by (c), we get $g_{2}^{-1}g_{1}g_{1}^{-1}x_{1}=g_{2}^{-1}x_{1}\in B_{10}(p)$. Hence by (b),
$d(x_{1},x_{2})$ $=d(g_{2}^{-1}x_{1},g_{2}^{-1}x_{2})$
$\leq d(f(g_{2}^{-1}x_{1}),f(g_{2}^{-1}x_{2}))+\delta$
$\leq d(\varphi(g_{2}^{-1}g_{1})f(g_{1}^{-1}x_{1}),f(g_{2}^{-1}x_{2}))+2\delta$
$\leq 5,$
so (3.5) applies, proving (2). For $y\in Y$, by (a) there is $g\in G_{X}$ with $d(\varphi(g)q,y)\leq\delta$, so
$d(\tilde{f}(gp),y)\leq d(\varphi(g)f(g^{-1}gp),y)+\delta\leq 2\delta,$
proving (3).
Now assume $Y$ satisfies (3.1), and $f$ is continuous. Then instead of defining $\tilde{f}$ by (3.4), define $\tilde{f}(x)\in Y$ to be the center of mass of $\mu_{x}$. As the functions $\rho_{g}$ are all continuous, $\mu_{x}$ depends continuously and equivariantly on $x$, so the result follows. ∎
When two geodesic spaces $X$ and $Y$ are Gromov–Hausdorff close, very little can be said about their topologies (see for example [FO95]). However, if $X$ satisfies nice local contractibility properties and admits a continuous Gromov–Hausdorff approximation $X\to Y$, some of its topological features are dominated by the corresponding ones of $Y$ (see for example [P90]). Lemmas 3 and 3 below are in the spirit of this phenomenon.
[…]
For each $m\in\mathbb{N}$, there is $\varepsilon>0$ such that the following holds. Let $X$, $Y$ be proper geodesic spaces and maps $f:X\to Y$, $h:Y\to X$ such that
  1. (1)
    For each $r\in(0,1]$, $x\in X$, the ball $B_{r}(x)$ is contractible in $B_{r+\varepsilon}(x)$.
  2. (2)
    For all $y_{1},y_{2}\in Y$ with $d(y_{1},y_{2})\leq 1$, one has $d(hy_{1},hy_{2})\leq d(y_{1},y_{2})+\varepsilon$.
  3. (3)
    For all $x\in X$, one has $d(h(f(x)),x)\leq\varepsilon$.
  4. (4)
    $f$ is continuous.
Then the map $f_{\ast}:H_{m}(X)\to H_{m}(Y)$ is injective. In particular, if $H_{m}(Y)$ is trivial, then so is $H_{m}(X)$.
Ideally, we would like the proof of Lemma 3 to go like this: $h\circ f\approx\text{Id}_{X}$, thus the identity map $H_{m}(X)\to H_{m}(X)$ factors as
$H_{m}(X)\xrightarrow{f_{\ast}}H_{m}(Y)\xrightarrow{h_{\ast}}H_{m}(X).$
This argument fails as $h$ is not necessarily continuous, so $h\circ f$ is not exactly homotopic to the identity on $X$. However, condition (1) allows us to get around these issues. We now turn to the actual proof.
Proof.
Assume $\varepsilon<\frac{1}{4m+3}$. For each $x_{0},\ldots,x_{k}\in X$ with ${\text{diam}}(\{x_{0},\ldots,x_{k}\})\leq 3\varepsilon$ and $k\leq m+1$, we will construct a singular $k$-simplex $[x_{0},\ldots,x_{k}]$ in $X$. For each $x_{0}\in X$, we denote by $[x_{0}]$ the singular 00-simplex consisting of the point $x_{0}$. For each $x_{0},x_{1}\in X$ with $d(x_{0},x_{1})\leq 3\varepsilon,$ let $[x_{0},x_{1}]$ be a geodesic from $x_{0}$ to $x_{1}$ seen as a singular 1-simplex.
Fix $k\leq m$ and assume by induction that for each $j\in\{1,\ldots,k\}$, and any points $x_{0},\ldots,x_{j}\in X$ with ${\text{diam}}(\{x_{0},\ldots,x_{j}\})\leq 3\varepsilon$, there is a singular $j$-simplex $[x_{0},\ldots,x_{j}]$ in $X$ contained in $B_{4j\varepsilon}(x_{0})$ and compatible with the lower dimensional simplices; that is,
(3.6) $\partial[x_{0},\ldots,x_{j}]=\sum_{\ell=0}^{j}\,(-1)^{\ell}\,[x_{0},\ldots,\hat{x}_{\ell},\ldots,x_{j}].$
Take $x_{0},\ldots,x_{k+1}\in X$ with ${\text{diam}}(\{x_{0},\ldots,x_{k+1}\})\leq 3\varepsilon$. Then the already constructed simplices $[x_{0},\ldots,\hat{x}_{j},\ldots,x_{k+1}]$ are contained in $B_{(4k+3)\varepsilon}(x_{0})$ and by (3.6) they satisfy
$\sum_{j=0}^{k+1}\,(-1)^{j}\,\partial\,[x_{0},\ldots,\hat{x}_{j},\ldots,x_{k+1}]=0$
so by (1) there is a singular $(k+1)$-simplex $[x_{0},\ldots,x_{k+1}]$ contained in $B_{4(k+1)\varepsilon}(x_{0})$ that restricted to its facets gives the ones already constructed; that is,
(3.7) $\partial[x_{0},\ldots,x_{k+1}]=\sum_{j=0}^{k+1}\,(-1)^{j}\,[x_{0},\ldots,\hat{x}_{j},\ldots,x_{k+1}].$
For each singular $k$-simplex $\tau:\Delta^{k}\to Y$, we define $h_{\sharp}(\tau)$ to be the singular $k$-simplex in $X$ given by $[h\tau(e_{0}),\ldots,h\tau(e_{k})]$, where $e_{0},\ldots,e_{k}$ are the vertices of $\Delta^{k}$. By (2) this can be done provided ${\text{diam}}(\tau(\Delta^{k}))\leq\varepsilon$ and $k\leq m+1$. This definition expands linearly to all chains of the form $\psi=\sum_{\tau\in T}n_{\tau}\tau$ where $T$ is a finite set of singular $k$-simplices of diameter $\leq\varepsilon$ and $n_{\tau}\in\mathbb{Z}$ for each $\tau$, as
$h_{\sharp}(\psi):=\sum_{\tau\in T}n_{\tau}h_{\sharp}(\tau).$
Notice that by (3.7), we also have $\partial h_{\sharp}(\psi)=h_{\sharp}(\partial\psi)$.
Consider an $m$-cycle $\varphi=\sum_{\sigma\in S}n_{\sigma}\sigma$ in $X$ representing a homology class in ${\text{Ker}}(f_{\ast})$, where $S$ is a finite set of singular $m$-simplices, and $n_{\sigma}\in\mathbb{Z}$ for each $\sigma\in S$. Our goal is to show that $\varphi$ is a boundary. By (4), the cycle
$f_{\sharp}(\varphi)=\sum_{\sigma\in S}n_{\sigma}f_{\sharp}(\sigma):=\sum_{\sigma\in S}n_{\sigma}(f\circ\sigma)$
is well defined and by hypothesis is a boundary. Take $\psi$ a singular $(m+1)$-chain in $Y$ with $\partial\psi=f_{\sharp}(\varphi)$. By applying barycentric subdivision simultaneously on both sides, we can assume the diameter of any singular simplex featured in the chains $\varphi$ and $\psi$ has diameter $\leq\varepsilon$, so $h_{\sharp}(\psi)$ is well defined.
Then by (3), the boundary
$\partial h_{\sharp}(\psi)=h_{\sharp}(\partial\psi)=h_{\sharp}f_{\sharp}(\varphi)=\sum_{\sigma\in S}n_{\sigma}h_{\sharp}f_{\sharp}(\sigma)$
is uniformly close to $\varphi$. That is, for each $\sigma\in S$, the maps $\sigma,\,h_{\sharp}f_{\sharp}(\sigma):\Delta^{m}\to X$ are at uniform distance $\leq 8m\varepsilon.$ Thus if $\varepsilon$ is small enough (depending on $m$) one can (inductively on the skeleta just like at the beginning of the proof) construct a homotopy between $\varphi$ and $h_{\sharp}f_{\sharp}(\varphi)$. This implies that $\varphi$ is also a boundary.
[…]
Let $X$, $Y$ be proper geodesic spaces and maps $f:X\to Y$, $h:Y\to X$ that satisfy conditions (1), (2), (3), (4) from Lemma 3 with $\varepsilon=1/100$. Then $f_{\ast}:\pi_{1}(X)\to\pi_{1}(Y)$ is injective. In particular, if $Y$ is simply connected then so is $X$.
Proof.
The proof is essentially the same as for Lemma 3 (c.f. [SW01, Theorem 2.1]). Given a loop $\varphi:\mathbb{S}^{1}\to X$ representing an element of ${\text{Ker}}(f_{\ast})$, there is a disk $\psi:\mathbb{D}\to Y$ with $\psi|_{\mathbb{S}^{1}}=f\circ\varphi$. Triangulate $\mathbb{D}$ in such a way that the image of each face of the triangulation under $\psi$ has diameter $\leq 1/100$. Then for each face with vertices $a,b,c\in\mathbb{D}$, construct the singular 2-simplex $[h(\psi(a)),h(\psi(b)),h(\psi(c))]$ in $X$ as in the proof of Lemma 3. Putting these simplices together, we obtain a continuous map $h_{\sharp}\psi:\mathbb{D}\to X$ so that $h_{\sharp}\psi|_{\mathbb{S}^{1}}$ is uniformly close to $\varphi$. Applying condition (1) once more we obtain that $\varphi$ is homotopic to $h_{\sharp}\psi|_{\mathbb{S}^{1}}$, hence nullhomotopic. ∎
[…]
At first, condition (4) in Lemma 3 may seem like an unnecessary technicality. However, simple examples such as Berger metrics on $\mathbb{S}^{3}$ converging to a round metric on $\mathbb{S}^{2}$ show that this condition is actually necessary.
[…]
With essentially the same proof, Condition (1) in lemmas 3 and 3 can be replaced by the existence of a contractibility function in the sense of [P90] (on which $\varepsilon$ depends). Since we will not need such version, we leave it as an exercise for the reader.
The following lemma is a basic result of algebraic topology.
[…]
Let $X$ be a simply connected topological manifold of dimension $n$. If $H_{m}(X)=0$ for all $m\in\{1,\ldots,n\}$, then $X$ is contractible.
Proof.
By [H02, Proposition 3.29] all homology groups of $X$ vanish, so by Hurewicz Theorem [H11, Theorem 2.1] and induction all homotopy groups vanish as well. By the work of Milnor [M59, Corollary 1], $X$ is homotopy equivalent to a CW-complex and Whitehead Theorem [H02, Theorem 4.5] applies, so $X$ is contractible. ∎
[…]
An infranilmanifold is defined to be the smooth manifold obtained as the quotient of a simply connected nilpotent Lie group $G$ by a torsion-free discrete subgroup $\Gamma$ of the semi-direct product of $G$ with a compact group of automorphisms of $G$.
[…]
Let $X$ be a closed aspherical manifold with $\pi_{1}(X)$ virtually nilpotent. Then $X$ is homeomorphic to an infranilmanifold.
The following argument is extracted from [K21].
Proof.
By the work of Lee–Raymond [LR82], there is an infranilmanifold $M$ with $\pi_{1}(M)=\pi_{1}(X)$. Then from the proof of the geometrization conjecture by Perelman in dimension $3$, the work of Freedman–Quinn in dimension $4$ [FQ90], and the work of Farrell–Hsiang in dimensions 5 and higher [FH83], the Borel conjecture holds for such groups, thus the manifolds $X$ and $M$ are homeomorphic. ∎
4. Asphericity Theorem
As mentioned in the introduction, a crucial step to prove Theorem 1 is to show that the space is aspherical. We now prove something slightly more general.
[…]
For each $K\in\mathbb{R}$, $N\geq 1$, there is $\varepsilon>0$ such that if $(X,\mathsf{d},\mathfrak{m})$ is an $\text{RCD}(K,N)$ space and $G\leq\text{Iso}(X)$ is a discrete group of isometries with ${\text{diam}}(X/G)\leq\varepsilon$, then
(4.1) ${\text{rank}}(G)\leq N,$
and in case of equality $X$ is homeomorphic to $\mathbb{R}^{N}$.
Proof.
Assuming the theorem fails, there is a sequence $(X_{i},\mathsf{d}_{i},\mathfrak{m}_{i})$ of $\text{RCD}(K,N)$ spaces and $G_{i}\leq\text{Iso}(X_{i})$ discrete groups of isometries with ${\text{diam}}(X_{i}/G_{i})\to 0$ such that for all $i$ either ${\text{rank}}(G_{i})>N$, or ${\text{rank}}(G_{i})=N$ and $X_{i}$ is not homeomorphic to $\mathbb{R}^{N}$. After taking a subsequence and re-scaling the spaces $X_{i}$ by factors $\lambda_{i}$ with $\lambda_{i}\to\infty$ and $\lambda_{i}\cdot{\text{diam}}(X_{i}/G_{i})\to 0$, we can assume $(X_{i},\mathsf{d}_{i},\mathfrak{m}_{i})$ is an $\text{RCD}(-\frac{1}{i},N)$ space for each $i$.
For $p_{i}\in X_{i}$, by taking again a subsequence, we can assume $(X_{i},p_{i})$ converges in the Gromov–Hausdorff sense to a proper geodesic space $(X,p)$. By Lemma [Z24, Theorem 4.1], $X$ is a nilpotent Lie group equipped with an invariant sub-Finsler metric. Since it is also an $\text{RCD}(0,N)$ space, it is Riemannian (see [J21, LDLP23]), and moreover by [M76, Theorem 2.4], abelian. By Lemma 2.8, there is $s\in\mathbb{N}$ and groups $G_{i}^{\prime}\leq G_{i}$ with (2.2) and (2.3).
By Lemma 2.8 and Remark 2.8, for $i$ large enough we have
(4.2) ${\text{rank}}(G_{i})={\text{rank}}(G_{i}^{\prime})\leq n\leq N,$
where $n$ denotes the topological dimension of $X$. This finishes the proof of (4.1).
From now on we assume ${\text{rank}}(G_{i})=N$ for all $i$, so equality holds in (4.2) and by Remark 2.8 we have $(X,p)=(\mathbb{R}^{N},0)$. Hence by Corollary 2.3, the groups $G_{i}^{\prime}$ admit no small subgroups. Then by Lemma 2.8 and Remark 2.8, for $i$ large enough there are elements $u_{1,i},\ldots,u_{n,i}\in\Gamma_{i}=G_{i}^{\prime}$, and $N_{1,i},\ldots,N_{n,i}\in\mathbb{R}^{+}$ with the following properties:
  • The set map $\alpha_{i}:\mathbb{Z}^{n}\to\Gamma_{i}$ given by
    $\alpha_{i}(m_{1},\ldots,m_{n})=u_{1,i}^{m_{1}}\cdots u_{n,i}^{m_{n}}$
    is a bijection. Moreover, there are polynomials $Q_{i}:\mathbb{R}^{n}\times\mathbb{R}^{n}\to\mathbb{R}^{n}$ of degree $\leq d(n)$ such that the product in $\Gamma_{i}$ is given by
    $g_{1}\cdot g_{2}=\alpha_{i}(Q_{i}(\alpha_{i}^{-1}(g_{1}),\alpha_{i}^{-1}(g_{2}))).$
  • There is $C>0$ such that the set
    $P_{i}:=P(u_{1,i},\ldots,u_{n,i};N_{1,i},\ldots,N_{n,i})\subset\Gamma_{i}$
    is a nilprogression in $C$-normal form with $\text{thick}(P_{i})\to\infty$.
  • For each $\varepsilon>0$ there is $\delta>0$ such that
    (4.3) $G(\delta P_{i})\subset$$\{g\in\Gamma_{i}\,|\,\mathsf{d}_{i}(gp_{i},p_{i})\leq\varepsilon\}$
    (4.4) $\{g\in\Gamma_{i}\,|\,\mathsf{d}_{i}$ $(gp_{i},p_{i})\leq\delta\}\subset G(\varepsilon P_{i})$
    for $i$ large enough.
By (4.4), after re-scaling the sequence $X_{i}$ by a fixed factor, we can assume
(4.5) $\{g\in\Gamma_{i}|\mathsf{d}_{i}(gp_{i},p_{i})\leq 100\}\subset G(P_{i}).$
By (4.3), there is $\delta_{1}>0$ with $G(\delta_{1}P_{i})\subset\{g\in\Gamma_{i}|\mathsf{d}_{i}(gp_{i},p_{i})\leq 1\}$. Hence there is $C_{1}\in\mathbb{N}$ so that
(4.6) $G(P_{i})\subset G(\delta_{1}P_{i})^{C_{1}}\subset\{g\in\Gamma_{i}|\mathsf{d}_{i}(gp_{i},p_{i})\leq C_{1}\}.$
Let $\mathfrak{l}_{i}$ be the Lie algebra of the Lie group $(\mathbb{R}^{n},Q_{i})$, and let
$\text{v}_{j,i}:=\log\left(u_{j,i}^{\lfloor\frac{N_{j,i}}{C}\rfloor}\right)\in\mathfrak{l}_{i}.$
By (4.6), after passing to a subsequence, for each $j\in\{1,\ldots,n\}$ we can assume
$u_{j,i}^{\lfloor\frac{N_{j,i}}{C}\rfloor}p_{i}\to\text{v}_{j}\,\,\text{ for some }\text{v}_{j}\in X=\mathbb{R}^{n}.$
By Lemmas 2.8 and 2.8, after further taking a subsequence, the groups $G_{i}^{\prime}$ converge equivariantly to the group of translations in $X=\mathbb{R}^{n}$. It is proven in [Z24, Lemma 8.2 and Lemma 2.64] that the structure coefficients of $\mathfrak{l}_{i}$ with respect to the basis $\{\text{v}_{1,i},\ldots,\text{v}_{n,i}\}$ converge to the structure coefficients in $\mathbb{R}^{n}$ with respect to $\{\text{v}_{1},\ldots,\text{v}_{n}\}$ as $i\to\infty$. These coefficients are of course 00 as $\mathbb{R}^{n}$ is abelian.
For each $i$, equip $\mathfrak{l}_{i}$ with the inner product $\langle\cdot,\cdot\rangle_{i}$ that makes the linear isomorphism $\mathfrak{l}_{i}\to\mathbb{R}^{n}$ given by $\text{v}_{j,i}\mapsto\text{v}_{j}$ an isometry and let $Y_{i}$ be the corresponding simply connected nilpotent Lie group equipped with the left invariant Riemannian metric induced from $\langle\cdot,\cdot\rangle_{i}$. Also let $q_{i}\in Y_{i}$ be the neutral element and consider the map
$f_{i}:(X_{i},p_{i})\to(Y_{i},q_{i})$
obtained by composing the Gromov–Hausdorff approximation $X_{i}\to X$ we have by definition, the isometric linear isomorphism $X\to\mathfrak{l}_{i}$ that sends $\text{v}_{j}\mapsto\text{v}_{j,i}$, and the exponential map $\mathfrak{l}_{i}\to Y_{i}$. Notice that by Theorem 2.4 we can assume $f_{i}|_{B_{100}(p_{i})}$ is continuous for $i$ large enough.
[…]
For each $\delta>0$, the following holds for $i$ large enough:
  1. (a)
    ${\text{diam}}(X_{i}/\Gamma_{i})$, ${\text{diam}}(Y_{i}/\Gamma_{i})\leq\delta$.
  2. (b)
    For all $x_{1},x_{2}\in B_{10}(p_{i})$,
    $|d(f_{i}(x_{1}),f_{i}(x_{2}))-\mathsf{d}_{i}(x_{1},x_{2})|\leq\delta.$
  3. (c)
    If $x\in B_{10}(p_{i})$, $g\in\Gamma_{i}$ are such that $gf_{i}(x)\in B_{9}(q_{i})$, then
    $gx\in B_{10}(p_{i}).$
  4. (d)
    If $x\in B_{10}(p_{i})$, $g\in\Gamma_{i}$ are such that $gx\in B_{10}(p_{i})$, then
    $d(f_{i}(gx),g(f_{i}(x)))\leq\delta.$
Proof.
(a) is satisfied by hypothesis and the fact that the structure coefficients of $\mathfrak{l}_{i}$ converge to 0. (b) is satisfied as the maps $X_{i}\to X\to\mathfrak{l}_{i}\to Y_{i}$ are each either an isometry or a map uniformly close to being an isometry on a large ball.
By contradiction, if (d) doesn’t hold, after passing to a subsequence we could find $x_{i}\in B_{10}(p_{i}),$ $g_{i}\in\Gamma_{i}$ such that $g_{i}x_{i}\in B_{10}(p_{i})$ and
(4.7) $d(f_{i}(g_{i}x_{i}),g_{i}f_{i}(x_{i}))\geq\delta\,\text{ for all }i.$
By (4.5),
$g_{i}=\exp(t_{1,i}\text{v}_{1,i})\cdots\exp(t_{n,i}\text{v}_{n,i})$
for some $t_{1,i},\ldots,t_{n,i}\in[-1,1]$. After passing again to a subsequence we have $t_{j,i}\to t_{j}$ as $i\to\infty$ for some $t_{j}\in[-1,1]$. Then by how we defined the vectors $\text{v}_{j}$, it follows that $\exp(t_{j,i}\text{v}_{j,i})$ converges to the translation in $X$ by the vector $t_{j}\text{v}_{j}$. As the convergence is equivariant, $g_{i}$ converges to the translation by $t_{1}\text{v}_{1}+\ldots+t_{n}\text{v}_{n}$.
On the other hand, by looking at the Gromov–Hausdorff approximations $Y_{i}\to X$ obtained by composing $\log:Y_{i}\to\mathfrak{l}_{i}$ with the linear isomorphism $\mathfrak{l}_{i}\to X$ that sends $\text{v}_{j,i}\mapsto\text{v}_{j}$, the sequence of actions of $\Gamma_{i}$ on $Y_{i}$ converge (again by Lemmas 2.8 and 2.8) to the group of translations of $X$. With respect to this convergence, by how we chose the map $\mathfrak{l}_{i}\to X$, the sequence $\exp(t_{j,i}\text{v}_{j,i})$ converges again to the translation in $X$ by $t_{j}\text{v}_{j}$, so $g_{i}$ again converges to $t_{1}\text{v}_{1}+\ldots+t_{n}\text{v}_{n}$.
The action of $g_{i}$ on both $X_{i}$ and $Y_{i}$ converges to the same isometry of $X$, and the Gromov–Hausdorff approximations $f_{i}:X_{i}\to Y_{i}$ were obtained by composing the ones we used for $X_{i}\to X$ with the approximate inverses of the ones we used for $Y_{i}\to X$, meaning that $f_{i}$ approximately commutes with the action of $g_{i}$, contradicting (4.7).
Lastly, assume (c) fails, so after taking a subsequence there are $x_{i}\in B_{10}(p_{i})$, $g_{i}\in\Gamma_{i}$ with $g_{i}f_{i}(x_{i})\in B_{9}(q_{i})$ but $g_{i}x_{i}\not\in B_{10}(p_{i})$ for all $i$. After taking again a subsequence both sequences $x_{i}$ and $f_{i}(x_{i})$ converge to a point $x\in X$ and $g_{i}$ converges to an isometry of $X$. On one hand, since $g_{i}f_{i}(x_{i})\in B_{9}(q_{i})$ for all $i$, we have $gx\in B_{19/2}(p)$, but on the other hand since $g_{i}x_{i}\not\in B_{10}(p_{i})$ for all $i$ we also have $gx\not\in B_{19/2}(p)$. This is a contradiction. ∎
Now fix $\varepsilon>0$ given by Lemma 3 with $m=n$. By Proposition 4, the conditions of Lemma 3 hold for $i$ large enough so we have continuous $\Gamma_{i}$-equivariant maps $\tilde{f}_{i}:X_{i}\to Y_{i}$ such that
  1. (1)
    For all $x_{1},x_{2}\in X_{i}$ with $\min\{\mathsf{d}_{i}(x_{1},x_{2}),d(\tilde{f}_{i}(x_{1}),\tilde{f}_{i}(x_{2}))\}\leq 2$ one has
    $|d(\tilde{f}_{i}(x_{1})\tilde{f}_{i}(x_{2}))-\mathsf{d}_{i}(x_{1},x_{2})|\leq\varepsilon/3.$
  2. (2)
    For all $y\in Y_{i}$, there is $x\in X_{i}$ with $d(\tilde{f}_{i}x,y)\leq\varepsilon/3$.
By (2) we can define $h_{i}:Y_{i}\to X_{i}$ so that $d(\tilde{f}_{i}h_{i}(y),y)\leq\varepsilon/3$ for all $y\in Y_{i}$. Then from (1) follows that
  • For all $y_{1},y_{2}\in Y_{i}$ with $d(y_{1},y_{2})\leq 1$, one has $\mathsf{d}_{i}(h_{i}y_{1},h_{i}y_{2})\leq d(y_{1},y_{2})+\varepsilon$.
  • For all $x\in X_{i}$, one has $\mathsf{d}_{i}(h_{i}(\tilde{f}_{i}(x)),x)\leq\varepsilon$.
By Corollary 2.4, for $i$ large enough
  • For each $r\in(0,1]$, $x\in X_{i}$, the ball $B_{r}(x)$ is contractible in $B_{r+\varepsilon}(x)$.
This means that all the conditions of Lemmas 3 and 3 are satisfied, hence the spaces $X_{i}$ are simply connected and have trivial homology up to degree $n$. By Corollary 2.4 the spaces $X_{i}$ are manifolds of dimension $n$ so from Lemma 3 they are contractible. Since the groups $\Gamma_{i}$ are discrete and torsion-free, they act freely and the quotient $X_{i}/\Gamma_{i}$ is a closed aspherical manifold. By Theorem 3, $X_{i}/\Gamma_{i}$ is homeomorphic to a nilmanifold, so its universal cover $X_{i}$ is homeomorphic to $\mathbb{R}^{n}=\mathbb{R}^{N}$. ∎
Proof of Theorem 1:.
By Theorem 4, there is $\varepsilon(K,N)>0$ such that if ${\text{diam}}(X)\leq\varepsilon$ for an $\text{RCD}(K,N)$ space $X$, then ${\text{rank}}(\pi_{1}(X))\leq N$, and in case of equality $X$ is an aspherical manifold. From the proof of Theorem 4 we deduce that $\pi_{1}(X)$ is virtually nilpotent, so the result follows from Theorem 3. ∎
5. Eigenmap to a flat torus
As pointed out in the introduction, the following result, which is a slight modification of [HP23, Theorem 6.3], finishes the proof of Theorem 1.1 (see also [H20, Corollary 1.9] for a more general statement valid for smooth manifolds).
[…]
Let $(X_{i},\mathsf{d}_{i},\mathfrak{m}_{i})$ be a sequence of compact $\text{RCD}(K,N)$ spaces that converges in the measured Gromov–Hausdorff sense to $\mathbb{T}^{N}$. Consider a sequence of abelian groups $G_{i}\leq\text{Iso}(X_{i})$ with ${\text{diam}}(X_{i}/G_{i})\to 0$. Then for $i$ large enough there is a free isometric action of $G_{i}$ on $\mathbb{T}^{N}$ and a $G_{i}$-equivariant homeomorphism $F_{i}:X_{i}\to\mathbb{T}^{N}$ with
$(1-\varepsilon_{i})\,\mathsf{d}_{i}(x,y)^{1+\varepsilon_{i}}\leq\mathsf{d}(F_{i}(x),F_{i}(y))\leq(1+\varepsilon_{i})\,\mathsf{d}_{i}(x,y)\,\text{ for all }x,y\in X_{i}$
for some $\varepsilon_{i}\to 0$. Furthermore, if the spaces $X_{i}$ are smooth Riemannian manifolds, the maps $F_{i}$ can be taken to be diffeomorphisms for $i$ large enough.
Before we jump into the proof, we make a few remarks. If $(X,\mathsf{d},\mathfrak{m})$ is an $\text{RCD}(K,N)$ space, and $G\leq\text{Iso}(X)$ a group of measure preserving isometries, then $G$ admits a natural isometric linear action on $H^{1,2}(X)$ by
$g^{\ast}f(x):=f(g^{-1}x),\,\,\,\,g\in G,f\in H^{1,2}(X),x\in X.$
This action clearly commutes with the Laplacian. That is, if $f\in H^{1,2}(X)$ is in the domain of the Laplacian, then $g^{\ast}f$ is also in the domain of the Laplacian, and
(5.1) $\Delta(g^{\ast}f)=g^{\ast}\Delta f.$
In particular, for each $\lambda\geq 0$, the eigenspace
$\{f\in H^{1,2}(X)|\Delta f=-\lambda f\}$
is a $G$-invariant subspace of $H^{1,2}(X)$.
Proof of Theorem 5:.
By Lemma 2.8, the sequence $G_{i}$ converges equivariantly to a connected abelian group $G\leq\text{Iso}(\mathbb{T}^{N})$ acting freely and transitively, which can be naturally identified with $\mathbb{T}^{N}$. It is well known that the spectrum of the Laplacian on $\mathbb{T}^{N}$ consists of non-positive integers, the multiplicity of the eigenvalue $-1$ is $2N$, and the eigenfunctions corresponding to that eigenvalue are given by the coordinate functions. That is, if $\varphi_{j},\psi_{i}:\mathbb{T}^{N}\to\mathbb{R}$ are given by
(5.2) $x=(\varphi_{1}(x),\psi_{1}(x),\ldots,\varphi_{N}(x),\psi_{N}(x))\,\text{ for all }x\in\mathbb{T}^{N},$
then
$\Delta\varphi_{j}=-\varphi_{j},\,\Delta\psi_{j}=-\psi_{j}\,\text{ for each }j\in\{1,\ldots,N\},$
and $\{\varphi_{j},\psi_{j}\}_{j=1}^{N}$ is a set of orthogonal functions in $H^{1,2}(\mathbb{T}^{N})$ with the same norm. Set
$W_{j}:=\langle\varphi_{j}\rangle\oplus\langle\psi_{j}\rangle=\langle G\varphi_{j}\rangle\hskip 85.35826ptV:=\bigoplus_{j=1}^{N}W_{j}\hskip 34.14322pt$
Then it is not hard to check that for $f\in V$, one has
[…]
For each $i\in\mathbb{N}$, set
$V_{i}:=\bigoplus_{1/2<\lambda<3/2}\{f\in H^{1,2}(X_{i})|\Delta f=-\lambda f\}$
By Corollary 2.5, for $i$ large enough, $\dim_{\mathbb{R}}(V_{i})=2N$, and by Lemma 2.7,
(5.4) $V_{i}=\bigoplus_{j=1}^{N}W_{j,i}$
with each $W_{j,i}$ 2-dimensional, $G_{i}$-invariant, and spanned by eigenfunctions.
[…]
Assume for some choice of $j(i)\in\{1,\ldots,N\}$, a sequence of functions $f_{i}\in W_{j(i),i}$ converges in $H^{1,2}$ to some function $f\in H^{1,2}(\mathbb{T}^{N})$. Then $f\in W_{j_{0}}$ for some $j_{0}\in\{1,\ldots,N\}$.
Proof.
By (5.3), if the lemma fails, there are $n_{k}\in\mathbb{N}$ for $k\in\{1,2,3\}$, $\alpha_{k,\ell}\in\mathbb{R}$, $g_{k,\ell}\in G$ for $\ell\in\{1,\ldots,n_{k}\}$, such that the vectors
$\text{v}_{k}:=\sum_{\ell=1}^{n_{k}}\alpha_{k,\ell}g_{k,\ell}^{\,\ast}(f)\in V$
with $k\in\{1,2,3\}$ form an orthonormal set. Take elements $g_{k,\ell,i}\in G_{i}$ with $g_{k,\ell,i}\to g_{k,\ell}$ as $i\to\infty$. Then for fixed $k$, the vectors
$\text{v}_{k,i}:=\sum_{\ell=1}^{n_{k}}\alpha_{k,\ell}g_{k,\ell,i}^{\,\ast}(f_{i})\in W_{j(i),i}$
converge to $\text{v}_{k}$ in $H^{1,2}$. Hence as $i\to\infty$, for $k\neq\ell$ one has
$\|\text{v}_{k,i}\|_{H^{1,2}}\to 1\hskip 85.35826pt\langle\text{v}_{k,i},\text{v}_{\ell,i}\rangle_{H^{1,2}}\to 0$
This is impossible as $\dim_{\mathbb{R}}(W_{j(i),i})=2$. ∎
By Theorem 2.5(2), for each $j\in\{1,\ldots,N\}$ there are eigenfunctions $\varphi_{j,i}\in V_{i}$ with $\varphi_{j,i}\to\varphi_{j}$ uniformly and in $H^{1,2}$.
[…]
For each $j\in\{1,\ldots,N\}$, the quantity
(5.5) $\min\{\|\varphi_{j,i}-\text{v}\|_{H^{1,2}}|\,\text{v}\in W_{\ell,i},\ell\in\{1,\ldots,N\}\}$
goes to 00 as $i\to\infty$.
Proof.
If the lemma fails, after passing to a subsequence, there is a sequence $\ell(i)\in\{1,\ldots,N\}$ for which we can write
$\varphi_{j,i}=\varphi_{j,i}^{a}+\varphi_{j,i}^{b}$
with $\varphi_{j,i}^{a}\in W_{\ell(i),i}$, $\varphi_{j,i}^{b}\in W_{\ell(i),i}^{\perp}\leq V_{i}$, and
$\min\left\{\liminf_{i\to\infty}\|\varphi_{j,i}^{a}\|_{H^{1,2}},\liminf_{i\to\infty}\|\varphi_{j,i}^{b}\|_{H^{1,2}}\right\}>0$
This is done by choosing $\ell(i)\in\{1,\ldots,N\}$ that minimizes (5.5), and defining $\varphi_{j,i}^{a}$ to be the orthogonal projection of $\varphi_{j,i}$ to $W_{\ell(i),i}$. Also let $\text{w}_{i}\in W_{\ell(i),i}$ be a unit vector orthogonal to $\varphi_{j,i}^{a}$. Since each $\varphi_{j,i}^{a}$, $\varphi_{j,i}^{b}$, $\text{w}_{i}$ is a linear combination of eigenfunctions in $V_{i}$, by Theorem 2.5(1) after passing to a subsequence we can assume that
$\varphi_{j,i}^{a}\to\text{v}^{a}\hskip 28.45274pt\varphi_{j,i}^{b}\to\text{v}^{b}\hskip 28.45274pt\text{w}_{i}\to\text{w}$
uniformly and in $H^{1,2}$ for some mutually orthogonal non-zero vectors $\text{v}^{a},\text{v}^{b},\text{w}\in V$. By Lemma 5, $\{\text{v}^{a},\text{w}\}$ forms a basis of $W_{j_{0}}$ for some $j_{0}$, and $\text{v}^{b}\in W_{j_{0}}^{\perp}\leq V$. This would mean that $\varphi_{j}=\text{v}^{a}+\text{v}^{b}+\text{w}$ is not in $W_{j}$; a contradiction. ∎
Let $\varphi_{j,i}^{a}$, $\text{w}_{i}$ be as in the proof of Lemma 5 and let $\psi_{j,i}^{a}:=\|\varphi_{j,i}^{a}\|_{H^{1,2}}\text{w}_{i}\in W_{\ell(i),i}$. Then
$\varphi_{j,i}^{a}\to\varphi_{j}\hskip 85.35826pt\psi_{j,i}^{a}\to\pm\psi_{j}$
uniformly and in $H^{1,2}$ as $i\to\infty$. After possibly multiplying $\psi_{j,i}^{a}$ by $-1$, we can assume $\psi_{j,i}^{a}\to\psi_{j}$. We point out that the awkwardness of $\varphi_{j,i}^{a}$ being in $W_{\ell,i}$ instead of $W_{j,i}$ emanates from the fact that we have a repeated eigenvalue in the limit space, and we have a priori no way of telling apart the spaces $W_{j,i}$ (see the discussion before [CC00, Theorem 7.3]).
Consider the map $\Phi_{i}:X_{i}\to V_{i}$ given by
$\Phi_{i}(x)=\sum_{j=1}^{N}[\varphi_{j,i}^{a}(x)\varphi_{j,i}^{a}+\psi_{j,i}^{a}(x)\psi_{j,i}^{a}]$
Fix $g\in G_{i}$. Since $W_{j,i}$ is $G_{i}$-invariant for each $j$, one has
$g^{\ast}\varphi_{j,i}^{a}$ $=\cos(\theta_{j,i})\varphi_{j,i}^{a}+\sin(\theta_{j,i})\psi_{j,i}^{a}$
$g^{\ast}\psi_{j,i}^{a}$ $=-\sin(\theta_{j,i})\varphi_{j,i}^{a}+\cos(\theta_{j,i})\psi_{j,i}^{a}$
for some angles $\theta_{j,i}\in\mathbb{S}^{1}$. Hence
$g^{\ast}\Phi_{i}(x)$ $=\sum_{j=1}^{N}[g^{\ast}\varphi_{j,i}^{a}(gx)g^{\ast}\varphi_{j,i}^{a}+g^{\ast}\psi_{j,i}^{a}(gx)g^{\ast}\psi_{j,i}^{a}]$
$=\sum_{j=1}^{N}[\cos^{2}(\theta_{j,i})+\sin^{2}(\theta_{j,i})][\varphi_{j,i}^{a}(gx)\varphi_{j,i}^{a}+\psi_{j,i}^{a}(gx)\psi_{j,i}^{a}]$
$=\Phi_{i}(gx),$
so $\Phi_{i}$ is $G_{i}$-equivariant. For each $i$, let $U_{i}\subset V_{i}$ denote the set of vectors whose projection to $W_{j,i}$ is non-zero for all $j$, and let $\pi_{i}:U_{i}\to U_{i}$ be the map given by
$\pi_{i}(z_{1},\ldots,z_{N}):=\left(\frac{z_{1}}{\|z_{1}\|_{H^{1,2}}},\ldots,\frac{z_{N}}{\|z_{N}\|_{H^{1,2}}}\right)$
with $z_{j}\in W_{j,i}$ for each $j$. Since the decomposition (5.4) is $G_{i}$-invariant, then so is the set $U_{i}$, and $\pi_{i}$ commutes with the action of $G_{i}$. Finally, we identify $V$ and $V_{i}$ with $\mathbb{R}^{2N}$ via the bases $\{\varphi_{1},\psi_{1},\ldots,\varphi_{N},\psi_{N}\}$ and $\{\varphi_{1,i}^{a},\psi_{1,i}^{a},\ldots,\varphi_{N,i}^{a},\psi_{N,i}^{a}\}$ respectively. After this identification, $\Phi_{i}$ converges uniformly to the isometric inclusion (5.2) of $\mathbb{T}^{N}$ in $\mathbb{R}^{2N}$. By Corollary 2.3 combined with Corollary 2.8, the action of $G_{i}$ on $X_{i}$ is free for large $i$. Then by Theorem 2.6 combined with Remark 2.6 and the fact that both $\Phi_{i}$ and $\pi_{i}$ are $G_{i}$-equivariant, the result follows. ∎